A Bibliography of Papers in Lecture Notes in Computer Science (2012): Volumes 7650–7699

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 1.01

Title word cross-reference

2 [187, 60, 15]. 3 [46]. ^{13} [332]. η_{T} [68]. GF(3^{97}) [68]. K [113, 108, 190]. λ [430]. N [481].

-Calculus [430]. -D [187]. -Means [190]. -Modes [113]. -Victims [481].

0/1 [325].

13th [498, 510, 503]. 15th [489]. 18th [490]. 19th [493, 494, 495, 496, 497].

2-Optimality [1]. 20th [492]. 256 [240].

3GPP [82]. 3GPP-MAC [82]. 3kf9 [82]. 3R [389].

4th [487, 488, 501].

51 [422].

8th [504, 500].

9th [502].

Abandoning [8]. ABE [307]. Abella [430]. ABML [109]. Abstract [353]. Accelerated [386]. Acceleration [212]. Accelerometer [446]. Accelerometer-Based [446]. Access [278, 276, 312, 306, 279, 322]. Accuracy [144]. Achieve [471]. Acquisition [253, 259]. Action [328, 361]. Activities [5, 134, 120, 205]. Activity [456, 457, 446, 463]. Ad [36, 98, 284]. Ad-Hoc [36, 284]. AdaBoost [7]. Adaptation [30]. Adaptive [386, 447, 457, 323, 453, 173, 328, 355]. Adaptively [73]. Adaptivity [83]. Additive [116]. Adjacencies [51]. Administration [312]. ADP [141]. Adult [257]. Advanced [440, 441]. Advances [503, 490]. Advantage [255]. Advertisers [371]. Aerial [204]. AES [244]. against [71, 485, 236, 285, 466, 80]. Agent [262, 453, 176, 334, 138]. Agglomerative [172]. Agreement [478, 367]. Agreement-Conflict [367]. Agricultural [404]. Aided [418]. AIGG [484]. Airport [318]. AIRS [504]. Alarm [118]. Alert [118]. Alertness [137]. Algebraic [75, 469]. Algorithm [388, 318, 391, 396, 321, 97, 323, 315, 395, 379, 187, 229, 333, 114, 173, 190, 105, 28, 394, 186, 385, 331, 384, 390, 185, 383, 45, 329, 212, 343, 269, 344]. Algorithms [50, 378, 491, 98, 177, 324, 52, 410, 61, 399]. Alias [420]. Alignment [27]. Allocation [197]. Always [126, 210]. Alzheimer [383]. Ambient [453, 461, 452, 509]. AmI [509]. Analaysis [489]. Analog [407]. Analyses [297]. Analysis [137, 262, 86, 332, 46, 270, 37, 239, 96, 491, 98, 119, 23, 408, 81, 78, 76, 136, 472, 291, 61, 420, 458, 384, 390, 392, 221, 300, 210, 178]. Analytics [400, 403, 399, 506]. Anchor [355]. Annealing [413, 168]. Anomaly [305, 299]. Anonymity [478]. Anonymity-Based [478]. Anonymous [301]. Answering [358]. Ant [333]. Antenna [382]. Anti [127, 105]. Anti-Histamine [127]. Anti-tampering [105]. Anycast [390]. Application [74, 332, 382, 177, 245, 189, 131, 387, 404, 168, 449, 490, 269]. Applications [73, 77, 29, 294, 139, 510, 440, 487, 488, 61, 30]. Applied [46]. Applying [29]. Approach [305, 142, 395, 67, 453, 153, 481, 261, 410, 105, 6, 339, 15, 306, 24, 441, 9, 12, 179, 220, 402, 373, 322, 141]. Approval [477]. Approximate [85]. Approximation [97, 229]. Approximations [115]. APT [481]. Arabic [157]. Architecture [303, 120, 455]. Area [252]. ARIA [472]. Arithmetic [191]. Array [382, 57, 355, 54]. Arrays [55]. Artificial [393, 396, 195, 105, 150, 208, 228]. ARX [78]. Asia [504]. ASIACRYPT [490]. ASMIA [300]. Aspects [296]. Assembler [422]. Assemblies [203]. Assessment [138, 275, 154, 384, 349]. Assignment [318]. Assistant [439]. Assistive [463]. Association [127, 108, 111, 106]. Associative [216, 252]. Assumptions [66, 242, 67]. Asymmetric [382]. Asymptotically [81].

ATESC [407]. Atmosphere [229]. Atomicity [290]. Attack [242, 305, 484, 240, 245]. Attacks [418, 467, 485, 241, 78, 285, 246, 80]. Attention [158]. Attentive [184]. Attitude [460]. Attribute [27]. Attributes [116]. Auction [289]. Audio [133, 337, 199]. Audio-Visual [133]. Audiovisual [337]. Auditory [137, 199]. Augmented [181]. August [510]. Austin [489]. Australia [501]. Authenticated [478]. Authentication [294, 476, 196, 306]. Auto [200]. Auto-encoder [200]. Automated [193, 203, 300]. Automatic [442, 451, 441, 247, 353]. Automation [418, 456]. Autonomous [287]. AutoProof [441]. Autoregressive [364]. Availability [301]. Aware [277]. Awareness [34, 453, 449]. Away [205]. Axiomatic [18]. Axiomatization [41].

Baby [428]. Back [103, 414, 443, 448]. Bacterial [383]. Bag [336]. Baggage [318]. Balance [225]. Balanced [471, 253]. Bands [35]. Base [14]. Based [58, 305, 243, 287, 396, 64, 447, 332, 290, 484, 270, 370, 11, 317, 204, 98, 119, 445, 446, 148, 407, 187, 68, 7, 229, 478, 324, 375, 397, 143, 226, 161, 225, 118, 105, 196, 283, 339, 27, 385, 369, 265, 360, 380, 15, 376, 390, 227, 8, 191, 316, 348, 151, 112, 12, 336, 149, 346, 208, 123, 349, 220, 201, 313, 304, 219, 373, 463, 322, 141, 180, 223, 345, 344, 250]. Baseline [342]. Bases [18]. Basic [50, 255]. Basis [342]. BatGaze [260]. Bayes [153, 366]. Bayesian [107, 184, 20, 213, 130, 158]. BB84 [236]. BDA [506]. Bear [450]. Bee [396]. Behavior [310, 451, 138, 408, 30, 201]. Behavior-Based [201]. Behavioral [461]. Behaviour [464]. Beijing [490]. Belief [18]. Benchmark [201]. Best [382]. Beta [170]. between [116, 267, 124, 253, 450]. beyond [82]. BFO [385]. Bhubaneswar [505]. BI [499]. Biasing [271]. Biclique [466]. Biclustering [226]. Bidding [138]. Big [403, 400, 506]. Big-Data [400]. Binary [49, 50, 62]. Binding [478]. Bio [166]. Bio-inspired [166]. Biochemical [330]. Birthday [82]. Blind [189]. Block [77, 350, 245, 76, 410, 472, 247, 178]. Boolean [469]. Bound [82]. Boundary [381]. Bounds [241]. BP [219]. Braille [448]. Brain [267, 127, 261, 263, 152, 499]. Breaking [68]. Breakout [327]. Bridge [124]. Broadside [382]. Browser [274, 479]. Browsing [406]. BTC [190]. Building [409, 312]. Bundling [340]. Business [120].

C [423, 332]. CA3 [165]. Cache [99]. Cache-Oblivious [99]. Caching [89]. Calculator [143]. Calculus [430]. Calibration [128]. Camellia [245]. Camera [352, 128]. Can [93]. Capacitive [446]. Car [448]. Case [46, 109, 214]. Cash [290]. Categorical [110]. Categories [145]. Categorization [166, 162]. Category [163, 358, 257]. Category-integrated [358]. Cauchy [338]. CBFO [385]. CBFO-Hybrid [385]. CBFO-S [385]. CEC2011 [202]. Cell [7]. Cellular [53]. Center [260]. Centric [475]. Centroid [168]. Certificate [292]. Certificate-Less [292]. Certificates [432, 306]. Certification [287, 419]. Certified [424, 507]. Change [89]. Channel [281]. Chaotic [161]. Character [200]. Characteristics [264].

Characterization [234]. Chattering [299]. Chattering-Free [299]. Checking [437, 419]. Chi [487, 488]. Children [450]. China [492, 504, 490, 499]. Chip [344]. Choice [3]. Chosen [244]. Chosen-Key [244]. Chronic [262]. Chunking [28]. Cipher [77, 81, 245, 76, 472, 80]. Ciphers [246, 247]. Circular [382]. Circularity [48]. City [487, 488]. Class [194]. Classification [401, 163, 207, 367, 479, 186, 366, 185, 228, 201, 368, 269]. Classifier [457, 324, 159]. Classifiers [298, 184, 409]. Click [360]. Client [483]. Client-Side [483]. Closest [333]. Closure [433]. Cloud [308, 403, 233, 309, 404]. CLPKM [292]. Clustered [100]. Clustering [393, 396, 395, 119, 407, 114, 173, 190, 172]. Clustering-Based [119]. Co [214]. Co-segmentation [214]. Code [424]. Coding [350, 341, 346]. Coercion [289]. Cognitive [134, 278, 135, 19, 205]. Coherently [187]. Coinductively [429]. Colaborative [31]. Collaboration [38]. Collaborative [310, 5, 2, 3, 361]. Collection [359]. Collective [23, 487, 488]. Collision [240]. Colony [396, 333]. Color [254, 168, 227]. Coloring [327]. Colorization [341]. Combination [144]. Combinatorial [60, 45, 489]. Combiner [17]. Combining [297, 366, 452]. Commerce [294]. Communication [86, 287, 450]. Community [358]. Commutation [57]. Compact [432]. Comparison [401, 1, 385, 17, 219, 361]. Compatible [199]. CompCert [423]. Compensation [350, 208]. Competition [202]. Competitive [216]. Compiler [417]. Compilers [421]. Completeness [433, 72, 234]. Complex [174]. Complexity [241, 136]. Components [49]. Composite [232]. Compositional [428]. Compositive [253]. Compression [91, 240]. Computation [94, 95, 237, 474, 143, 179]. Computational [164, 199, 487, 488]. Compute [231]. Computer [418, 191]. Computer-Aided [418]. Computing [249, 77, 29, 403, 505, 309, 404, 421]. Concepts [124]. Concerns [176]. Concurrent [426, 316]. Conditionals [14]. Conditions [411, 138]. Conference [502, 491, 498, 507, 504, 493, 494, 495, 496, 497, 503, 487, 488, 505, 509, 506, 500, 490, 499]. Conflict [367, 3]. Connected [49]. Connectivity [130]. Consensus [3]. Consensuses [1]. Consistency [308]. Constant [66, 93]. Constant-Size [66]. Constrained [341, 318]. Constraints [60]. Construction [298, 32]. Constructions [66, 74, 78]. Constructive [433]. Consumption [471, 150, 345]. Contact [342]. Content [337, 462, 301]. Contents [412]. Context [34, 445, 453, 462]. Context-Based [445]. Contextual [54, 459]. Contribution [258]. Control [98, 297, 451, 278, 276, 312, 279, 258, 141]. Controlled [473]. Controller [216]. Convergence [381]. Conversational [143]. convex [139]. Convolution [332, 59]. Cooperative [332]. Coordinates [24]. Coq [439]. Core [37, 316]. Correction [4, 342]. Correctness [422]. Correlation [79, 187]. Cortex [134, 205]. Cortical [137]. Costs [86]. Counting [221]. Coupled [222]. Covariance [181]. Cover [97]. Covers [44]. Covert [281]. CPP [507]. CPUs [316]. Creating [449]. Creativity [264]. Credit [155]. Crisis [451]. criteria [154]. Criterion [1]. Crop [404]. Cross [187].

Cross-Correlation [187]. Crossover [202]. Crowdsourcing [20]. Crude [15]. Cryptanalysis [243, 466]. Cryptographic [291]. Cryptography [418, 64, 470]. Cryptology [498, 490]. Cryptosystems [68]. CSS [501]. CT [218]. Cube [430]. Cubes [23]. Cubing [23]. Cuckoo [388]. Cultural [391]. Culture [165]. Current [471]. Curvatures [46]. Curve [47]. Curves [474]. Customer [192]. Cyberspace [501]. Cylindrical [24].

D [46, 187, 60]. Dafny [442]. Dam [189]. Darknet [310]. Data [116, 107, 270, 115, 155, 403, 26, 173, 225, 28, 136, 106, 387, 409, 47, 506, 295, 24, 231, 130, 400]. Database [308, 25, 288]. Databases [403, 13, 402, 405]. Datasets [110]. DC [386]. DCEL [44]. Dead [141]. Dead-Zone [141]. Deadline [101]. DEAL [329]. DEBORA [123]. December [502, 492, 491, 498, 507, 504, 503, 505, 506, 500, 490, 501, 499]. Decision [152, 39, 117, 154]. Decision-Making [152]. Decisional [70]. Decoding [267, 134]. Decomposition [178]. Decoupled [187]. Dedicated [147]. Deduplication [28]. Defence [285]. Definition [23]. Deflection [37]. Dehazing [229]. Delay [390]. Delayed [174]. Delhi [506]. Demand [292]. Dementia [464]. Dense [110]. Density [322]. Dependency [331, 123]. Dependency-Based [123]. Deployment [33]. Depth [261, 260, 352]. Design [129, 39, 324, 203, 322, 491]. Designs [418]. Detecting [85, 482, 370, 297, 483, 268, 300]. Detection [298, 310, 305, 401, 372, 484, 11, 299, 317, 445, 7, 206, 118, 339, 191, 151, 223]. Determination [207]. Determine [233, 481]. Determining [130]. Development [143, 164]. Deviations [217]. Device [268, 211]. Devices [296, 295]. Diabetic [193]. Diagnosis [383]. Dictionaries [99]. Dictionary [198]. Difference [216]. Differential [242, 382, 467, 215, 78, 76, 394, 472, 326, 380, 316, 80]. Differentials [247]. Digital [290, 46, 43, 283, 51, 306]. Digitized [228]. Dilution [135]. Dimensional [218]. Direction [329]. Direction-Guided [329]. Dirichlet [171, 197]. Disambiguation [434, 363]. Discovering [122, 8]. Discovery [323, 373]. Discrete [249, 379, 174, 60, 47]. Discriminant [213]. Disease [383]. Disorders [164]. Dispatch [385]. Dispersed [321]. Display [452]. Displays [462]. Disruption [390]. Dissection [232]. Dissimilarity [140]. Distant [355]. Distinguishers [79, 244]. Distortion [346]. Distributed [95, 187, 324, 394, 304]. Distribution [129, 394, 322]. Distributions [171]. Diversification [397]. Diversified [359]. Diversity [217, 159]. Division [435]. DOA [187]. Document [169, 197, 376]. Documentation [405]. Does [126, 210]. Doha [493, 494, 495, 496, 497]. Domain [475, 282, 106, 347, 123]. Dominance [319]. dominated [330]. Dominating [61]. Dominators [421]. Double [246]. Double-SP [246]. Down [101, 369]. Driven [462, 120, 455]. DRM [309]. Drugs [127, 300]. Dual [67, 181]. Duration [337]. during [255, 33, 256, 260, 136]. DyCML [471]. Dynamic [396, 95, 379, 100, 297, 33, 166, 138, 471, 173, 28, 276, 130]. Dynamical [194]. Dynamics [262, 305, 266, 381].

E-exam [196]. Early [152, 8, 383]. Eavesdropping [236]. Economic [385, 392, 408]. Economics [408]. EcoSim [320]. Edge [317]. Editor [31]. Education [404]. EEG [271, 136, 268]. Effect [264, 320, 215, 261, 208]. Effective [130]. Effects [135]. Efficiency [350]. Efficient [388, 74, 169, 89, 91, 482, 232, 323, 100, 25, 333, 412, 410, 409, 468, 269]. Effort [393]. Efforts [176]. EFP [25]. EFP-M2 [25]. Elapsed [90]. Elastic [163]. Elba [508]. Electronic [289]. Elements [203]. Elicitation [402]. Elimination [435]. Email [298, 483]. Embedded [435, 470, 250]. Embodied [450]. Emergence [320]. Emerging [33]. Emotion [268]. Emotional [119]. Emotive [268]. Emphasis [264]. Empirical [153, 210]. Employees [196]. Empowerment [264]. encoder [200]. Encrypted [231]. Encryption [232, 307, 304, 309]. End [346]. End-to-End [346]. Energetic [282]. Energy [87, 89, 150, 392]. Engineering [400, 379]. Enhanced [372, 402]. Enhancement [388]. Enhancing [95, 446, 375, 404, 360, 82]. Enough [365]. Enrichment [353]. Ensemble [298, 11, 324, 217, 409, 159]. Entailment [407]. Entity [365, 9, 123]. Entity-Relationship [123]. Entrepreneurial [126, 210]. Entropy [380]. Enumeration [469]. Environment [396, 458, 301, 208]. Environments [447, 460, 454, 279, 455]. Epoc [268]. ERP [255]. Errors [434]. Estimate [217]. Estimating [153, 128]. Estimation [129, 187, 195, 381, 348, 258, 199]. Evaluation [337, 234, 360, 452, 468, 383, 313]. Even [81, 469]. Even-Mansour [81]. Even-Variable [469]. Event [119, 120, 339, 455]. Event-Driven [120, 455]. Evoked [137]. Evolution [502, 382, 394, 326, 380, 334, 316]. Evolutionary [318, 315, 325, 505, 461, 329, 179]. Evolved [330]. Evolving [170]. exam [196]. Examinations [207]. Executable [423]. Execution [33]. Exemplar [351]. Expansion [86, 365]. Experimental [96]. Experts [4]. Explain [261]. Explicit [452]. Exploitation [114]. Exploiting [374, 405, 368, 357]. Exploration [386, 114]. Exploratory [361]. Exploring [374, 450, 357]. Explosion [437]. Extended [337, 11]. Extensive [117]. Extracting [123]. Extraction [386, 142, 218, 10, 362, 461, 9]. Extreme [36, 119]. Eye [342].

Face [214, 336]. Faceted [406]. Facial [341, 205, 336]. Factors [150, 313]. Factory [452]. Failure [99]. Fairness [176]. Fall [445, 191]. Family [242]. Fast [86, 59, 44, 330, 45, 178]. Faster [249, 94, 244]. Fatigue [262]. Fault [242, 92, 472]. Fault-Tolerant [92]. FBONT [170]. FCSRs [243]. Fear [408]. Feasibility [392]. Feasible [315]. Feature [133, 372, 270, 171, 218, 339, 340, 461, 192]. Features [375, 408, 260, 441, 228]. Feedback [206, 258]. Feistel [246]. Fermat [431]. Fiat [235]. FIFO [102]. Fighting [273]. Figure [317]. File [301]. File-Sharing [301]. Filling [343]. Filter [169, 118]. Filtering [298, 482]. Filters [129]. Financial [154]. Finding [107, 43]. Fingerprint [313]. FIREFLY [385]. Firing [153]. First [90, 323, 491, 506]. First-Order [323]. Fitness [317, 320]. Fitting [47]. Fixed [191]. Fixed-Point [191]. FJSSP [379]. Flexible [170, 216]. Flood [384]. Flooding [484, 485]. Flow [88, 297].

Flow-Time [88]. Flux [332]. FMRI [130, 270, 252, 256]. Focus [192]. Foraging [383]. Forecasting [170, 144, 138, 408]. Forecasts [35, 144]. Forgery [483]. Form [67, 117]. Formal [416, 435]. Formalization [430]. Formalizing [421]. Formally [420]. Formally-Verified [420]. Foundations [276, 492]. Fractional [131]. Fractional-Space [131]. Framework [49, 29, 16, 278, 355]. Free [299, 260]. Frequent [25, 12]. Front [21, 42, 63, 84, 104, 125, 146, 167, 188, 209, 230, 251, 272, 293, 314, 335, 356, 377, 398, 415, 436, 444, 465, 486]. Frontal [205]. Fuel [150]. Full [478, 475]. Function [240, 234, 342]. Functional [393, 424]. Functionalities [72]. Functionalities- [72]. Functionality [201]. Functions [74, 480, 317, 213, 241, 75, 469]. Fundus [193, 151]. Fusion [133, 184, 219, 180]. Future [64, 13]. Fuzzy [393, 266, 324, 16, 160, 283, 19, 15, 454, 172, 17, 194].

GA [330]. Gait [345, 344]. Games [117]. Gaps [94]. Garbling [73]. Gaussian [148, 151, 130]. Gaze [260]. GBML [324]. Gedi [491]. Gene [321]. General [307]. Generalized [74, 171, 241, 474, 381, 172, 194]. Generation [394, 463]. Generational [316]. Generator [100]. Generators [243, 291]. Generic [66, 22]. Genes [320]. Genetic [378, 317, 177, 114, 384, 390]. Geo [369]. Geo-locating [369]. Geographical [10]. GESNIC [485]. Gestural [455]. Gesture [352, 455]. GET [484, 485]. Global [194]. Gold [408]. Good [365]. GPU [226]. GPU-Based [226]. GPUs [468]. Gradients [7, 215]. Grading [193]. Grain [242]. Grammars [56]. Granular [19]. Granularity [177]. Graph [86, 114, 338, 269]. Graphics [212]. Graphs [61]. Gravitational [321]. Gray [388, 341, 190]. Greenhouse [208]. Grid [53, 302, 322, 302]. Group [181, 462, 387]. Growing [132]. Guided [326, 329]. Guwahati [500].

H.264 [347, 346]. H.264/AVC [347]. H.264/SVC [347]. Hamsi [240]. Hamsi-256 [240]. Hand [352]. Handling [387]. Handwriting [156]. Handwritten [157]. Hanoi [502]. Hard [71, 107]. Hard-to-Invert [71]. Harmony [202, 379]. Hash [480, 241, 475]. Hashing [338, 250]. Hastings [186]. Head [371]. Hearing [137]. Heartbeat [401]. Hebbian [164]. HEMH2 [325]. Hessenberg [410]. Hessian [474]. Heterogeneous [248, 102]. Heuristic [256, 2, 159, 259]. Hexagonal [56]. HID [305]. Hierarchical [374, 2]. High [350, 248]. High-Speed [248]. Higher [419, 245]. Higher-Order [419, 245]. Hilbert [194]. Hippocampal [165]. Histamine [127]. Histogram [7]. History [370]. Hoc [36, 98, 284]. HOG [223]. Hold’em [40]. Hole [343]. Hole-Filling [343]. Holistic [253]. Home [456, 458]. Homepage [373]. Hopfield [161]. Host [311]. HTTP [484, 485]. Human [119, 7, 365, 179, 223]. Human-Input [365]. Human-Simulated [179]. Hybrid [91, 391, 325, 385, 331, 156, 347, 392, 9]. Hybrid-Domain [347]. Hypergraph [263]. Hypergraphs [97]. Hyperplanes [45]. hypothesis [350].

I/O [410]. I2P [301]. ICCCI [487, 488]. ICISS [500]. Iconic [296]. ICONIP [493, 494, 495, 496, 497]. Identification [133, 321, 331, 149]. Identifiers [376]. Identity [294, 304]. Identity-Based [304]. IEEE [202]. IEEE-CEC2011 [202]. II [494, 488, 210]. III [495]. Illicit [300]. Im [69]. Image [388, 58, 489, 224, 62, 229, 375, 341, 351, 61, 168, 380, 384, 227, 149, 349, 219, 128]. Image-Based [375]. Images [193, 49, 50, 169, 140, 204, 190, 283, 151, 313]. Imaging [220]. Imbalanced [155]. Immune [75, 179]. Immunity [469]. Implement [206]. Implementation [485, 178]. Implementations [248]. Implicit [452]. Impossible [467, 247]. Improved [467, 325, 26, 190, 472, 392, 149, 349, 220]. Improving [155, 301]. Impulse [169]. Impulsive [174]. IMU [345, 344]. IMU-Based [345, 344]. Incoherent [236]. Incomplete [115]. Incorporating [375]. Incremental [58, 162, 112, 336]. Independent [129, 281]. Index [41, 13, 8]. India [498, 505, 506, 500]. Indirect [150]. Individual [264, 132, 361]. INDOCRYPT [498]. Induction [112]. Inductive [424, 257]. Inertial [445]. Infer [461]. Inference [283]. Influence [315, 254]. Informatics [261, 499]. Information [270, 255, 62, 504, 493, 494, 495, 496, 497, 375, 252, 503, 226, 367, 10, 265, 360, 312, 183, 490, 402, 357, 504, 510, 500]. Information-Theoretic [270]. Informative [370]. Initial [315]. Injection [348]. Input [107, 225, 365, 141]. Insider [288]. inspired [166]. Instance [175]. Instrument [206]. Integral [79]. Integrated [480, 343, 358]. Integrating [4]. Integrity [288]. Intel [422]. Intellectual [480]. Intelligence [105, 487, 488, 509, 461, 300]. Intelligent [492, 203, 118, 454]. Intentions [126, 210]. Intents [280]. Interaction [460, 448]. Interactive [177, 203, 4, 464]. Interface [39]. Interfaces [447, 455, 463]. International [489, 502, 492, 498, 507, 493, 494, 495, 496, 497, 510, 508, 487, 488, 505, 509, 506, 500, 490, 501, 499]. Interview [207]. Intrinsic [264, 334]. Introduction [438, 439, 330]. Intrusion [118]. Invariance [60]. Inverse [181]. Invert [71]. Investigation [378]. Investigations [481]. Investor [408]. Involving [194]. iOS [295]. IPv6 [476]. Isabelle [425]. Island [510, 508]. ISMIS [492]. iSOM [162]. Isothetic [43]. Isotopic [332]. Israel [491]. Italy [508, 509]. Itemset [110, 111]. Itemsets [12]. Iterate [74]. Iterated [241, 81]. Iterative [141, 194]. IV [496]. IWCIA [489].

Japan [507]. Jeju [510]. Joint [509, 219]. Juggle [429].

Kernel [157, 387]. Kernels [427]. Key [286, 244, 478, 476, 233, 292, 470, 304]. Keyer [448]. Keyless [282]. Keyphrase [362]. Keypoint [214]. Keystroke [305]. Kibbutz [491]. Knapsack [325]. Knapsacks [232]. Knowledge [142, 109, 5, 118, 3, 365, 106, 14, 363]. Knowledge-Based [118]. Kolkata [498]. Korea [510]. KP [307]. KP-ABE [307]. Kyoto [507].

Label [257]. Labeling [49, 332]. Labs [404]. Lagrangian [181]. Language [121, 177, 358, 253]. Languages [447, 57, 54]. Large [331, 402]. LASER [508]. Latency [77]. Latent [197]. Lattices [65]. Law [165]. Layer [124]. Layers [112]. Layout [322]. LBlock [466]. LBP [223]. Leakage [238, 71, 237]. Learn [225]. Learning [58, 401, 502, 457, 184, 7, 213, 139, 397, 182, 256, 161, 164, 189, 15, 454, 183, 9, 336, 159, 175, 395]. Lectures [508]. LED [76]. Less [292]. Level [388, 255, 59, 1, 473]. Level-Wise [59]. Leveraging [351, 365]. Lexical [124, 363]. License [222]. Lightweight [483]. Like [161]. Likelihood [185]. Limb [258]. Limen [215]. Line [157, 456, 237]. Line/Off [237]. Linear [79, 432, 341, 15, 169]. Link [393, 33]. Linked [178]. Linking [222]. List [287]. List-Based [287]. Little [431]. Load [144, 385]. Local [327, 115, 140, 375, 397]. Locality [341]. Locality-Constrained [341]. Localization [222]. Localized [171]. locating [369]. Location [382, 445]. Log [479]. Logarithms [249]. Logic [456, 432, 433, 438, 471, 160, 15]. Logical [106]. Long [337]. Look [236]. Loop [109]. Low [77]. Low-Latency [77]. LS [148]. LS-SVR [148]. Łukasiewicz [160]. Luminance [215]. Luring [411].

M [294]. M-Identity [294]. M2 [25]. MAC [82]. Macau [492, 499]. Machine [401, 267, 203, 139, 416]. Machines [11, 182, 147]. Macro [408]. Macro-economic [408]. Macula [151]. Maculopathy [193]. Mahalanobis [213]. Make [93]. Making [152, 154]. Malicious [273]. Malicous [482]. malleability [235]. Mammograms [228]. Management [476, 292, 284, 192, 304]. Manager [428]. MANET [278]. MANETs [292]. Manifold [139]. Manipulator [389]. Mansour [81]. Map [162]. Mapping [26, 328]. Mappings [194]. MapReduce [29]. Maps [207, 10, 19, 186, 145, 132]. Markerless [352]. Market [138]. Marketing [176]. Markov [456]. Mass [332, 321, 228]. Mass-Dispersed [321]. Matching [350, 214, 223]. Materialized [413]. Mathematical [411, 65]. Matrix [86, 224, 187]. Matter [103, 414, 443, 21, 42, 63, 84, 104, 125, 146, 167, 188, 209, 230, 251, 272, 293, 314, 335, 356, 377, 398, 415, 436, 444, 465, 486]. Max [110]. Maximum [469]. MCS [422]. MCS-51 [422]. Means [35, 190]. Measure [48, 111, 260]. Measurement [215, 345, 344]. Measurements [332]. Measures [217]. Measuring [274]. Mechanized [417, 421]. MedAlg [491]. Media [119, 406, 300]. Mediterranean [491]. Meet [245]. Meet-in-the-Middle [245]. Meeting [463]. Melbourne [501]. Melodic [206]. Memetic [142, 323, 505, 331]. Memory [89, 213, 428]. Merging [18]. Merit [317]. MESFET [386]. Mesh [285]. Messages [370]. Meta [395]. Meta-learning [395]. Metabolic [332]. Metadata [284]. Metaheuristics [325]. Meter [303]. Method [35, 372, 482, 204, 148, 181, 248, 26, 2, 387, 409, 15, 112, 149, 123, 180, 194]. Methods [401, 483, 154, 452, 17]. Metropolis [186]. Microprocessor [422]. Microprocessors [470]. Middle [245]. Min [110]. Min-Max [110]. Minh [487, 488]. Minimization [87]. Minimum [88, 327]. Mining [116, 108, 25, 111, 6, 106, 12]. Missing [195]. MISTY1 [467]. Mitigating [288]. MIX [289]. Mixture [171, 151]. Mobile [286, 274, 296, 294, 476, 284, 449, 211]. Modal [433, 14, 323]. Mode [299, 471]. Model [34, 386, 58, 262, 447, 121, 37, 437, 321, 25, 229, 166, 358, 419, 475, 364, 351, 278, 160, 164, 404, 165, 54, 316, 151, 279, 18, 263, 126, 199, 357]. Model-Based [447]. Modeling [40, 5, 135, 265, 15, 346, 158]. Modelling [411, 150, 156]. Models [416, 154, 312]. Modes [113]. Modifications [366]. Modified [148, 389]. Modular [195]. Moments [217]. Money [290]. Monitoring [189]. Morphological [61]. Motifs [8]. Motion [350, 339]. Motivating [176]. Motivation [264]. Motives [334]. Movement [408]. Movement-Forecasting [408]. Moving [13]. MRI [218]. Multi [298, 350, 270, 457, 202, 323, 453, 389, 176, 154, 384, 470, 334, 183, 316, 473, 279, 178, 180]. Multi-Agent [453, 176, 334]. Multi-block [178]. Multi-Classifier [457]. Multi-core [316]. Multi-criteria [154]. Multi-hypothesis [350]. Multi-Level [473]. Multi-modal [323]. Multi-objective [389]. Multi-Organization [279]. Multi-Parent [202]. Multi-precision [470]. Multi-sensor [384, 180]. Multi-Task [183]. Multi-tier [298]. Multi-Voxel [270]. Multicast [87]. Multiclass [147, 185]. Multicollision [241]. Multidimensional [79, 26, 8]. Multilevel [380]. Multimaps [99]. Multimedia [503, 339]. Multimodal [460]. Multinomial [366]. Multiobjective [325, 319]. Multipath [92, 208]. Multiple [121, 232, 184, 7, 182, 4, 313, 175]. Multiple-Instance [175]. Multiplication [271, 86, 470]. Multiprocessor [90]. Multivariable [24]. Multivariate [468]. Music [375, 374, 22]. Mutation [382, 378, 330]. Mysteries [65].

Naive [366, 263]. Name [363]. Named [9]. Nearly [141]. Necessity [411]. Needs [16]. Negligible [99]. Negotiation [117, 18]. Net [163]. Nets [216, 55]. Network [107, 36, 456, 321, 184, 213, 118, 189, 168, 191, 281, 208, 228, 130, 126, 210, 201, 222]. Network- [36]. Networking [3]. Networks [286, 393, 95, 200, 96, 204, 98, 100, 33, 195, 174, 161, 285, 284, 176, 369, 150, 390, 263]. Neural [393, 170, 224, 204, 213, 195, 134, 153, 226, 161, 189, 205, 150, 168, 191, 208, 228, 222, 493, 494, 495, 496, 497]. Niche [384]. Niching [397]. Node.js [275]. Nodes [37, 213]. Noise [129, 169, 348]. Noisy [319, 47]. Nominal [116]. Non [169, 456, 235, 108, 139, 475, 319, 330, 355, 450]. Non-convex [139]. Non-dominated [330]. Non-linear [169]. Non-malleability [235]. Non-parametric [319]. Non-programmable [475]. Non-redundant [108]. Non-verbal [450]. Non-visual [456]. Noninterference [427, 426]. Nonlinear [141, 128]. Nonparametric [171]. Normalization [121]. Normalized [157]. Novel [286, 143, 173, 13, 61, 306, 304]. November [489, 493, 494, 495, 496, 497, 487, 488, 509]. NP [107]. NP-Hard [107]. Number [315, 222].

O [410]. Object [46, 11, 43, 198, 254, 13, 214, 441, 402]. Object-Oriented [441]. objective [389]. Objects [162]. Oblique [261]. Oblivious [99]. OBS [37]. Observer [299]. Observing [310]. Obtaining [197]. OCE [31]. Off [157]. Off-Line [157]. Office [196]. Offline [200]. Oil [15]. OMP [219]. On-Chip [344]. On-Demand [292]. On-Line [456, 237]. On-Line/Off-Line [237]. One [74, 73, 1, 360]. One-Level [1]. One-Time [73]. One-Way [74]. Online [31, 129, 101, 143, 336]. Only [382]. Ontology [27, 124]. Open [123]. Open-Domain [123]. Operating [427]. Operations [22]. Operator [170, 319]. Operators [114]. Ophthalmic [207]. Opponent [40]. Optimal [236, 394, 70, 141]. Optimality [1]. Optimising [422]. Optimization [391, 332, 315, 333, 139, 381, 147, 160, 389, 319, 265, 331, 191, 392, 383, 322]. Optimizations [330]. Optimized [253]. Optimizer [386]. Options [38]. Oracle [475]. Orange [354]. orBAC [279]. Order [238, 323, 419, 245, 69]. Ordering [38]. Organising [186]. Organization [279]. Organized [218, 292]. Organizing [207, 367, 162, 145, 132]. Oriented [7, 362, 347, 441, 247]. Orthogonal [44]. Outlooks [182]. Outsourcing [73]. Over-constrained [318]. Overcomplete [198]. Overlapping [271]. Overview [10]. Oz [459].

P [54]. Pacific [503]. Pacific-Rim [503]. Packets [102]. Pages [482]. Pairing [64, 68, 474]. Pairing-Based [64, 68]. Palm [149]. Panel [410]. Papers [510]. PARAFAC [178]. PARAFAC/CP [178]. Parahippocampal [252]. Parallel [324, 248, 468]. Parallelism [152]. Parallels [267]. Parameter [386, 321, 378]. Parameters [128]. parametric [319]. Parent [202]. Parents [450]. Part [493, 494, 495, 496, 497, 487, 488, 223, 126, 210]. Part-Templates [223]. Partial [97, 57, 192]. Particle [386, 396, 381, 392]. Party [72]. Passwords [296]. Past [64]. Patch [140, 336]. Patch-Based [336]. Path [43]. Paths [95]. Pattern [270, 254, 6, 185, 9]. Patterns [85, 332, 94, 25, 461, 463]. Payment [477, 290]. PayWord [477]. PCM [503]. PDF [129]. PEAQ [199]. Pearl [430]. Pearls [431]. Pedestrian [221]. Peer [29]. Peer-to-Peer [29]. Penalized [185]. 1 [325]. AVC [347]. CP [178]. Off-Line [237]. SVC [347]. Perception [265]. Perception-Based [265]. Perceptual [152]. Perfect [75]. Performance [35, 37, 315, 364, 217, 383]. Periodic [85]. Periods [337]. Peripheral [254]. Permissions [449]. PermissionWatcher [449]. Person [262]. Personalization [462]. Personalized [32]. Perspective [400, 403]. Persuasive [452]. Perturbation [116]. Pervasive [77]. PET [220]. Petri [55]. Phase [89, 333, 471]. Phase-Change [89]. Phishing [298]. Phrases [122]. Physical [480, 312]. Physiological [266]. Piles [364]. Pisa [509]. Pitch [206]. Pixel [11, 60]. Pixel-Based [11]. Pixel-Invariance [60]. Pixels [282]. Place [252]. Placement [394]. Planar [128]. Plane [51]. Planning [91]. Plate [222]. Platform [23, 275]. Play [132]. Player [206]. PLIDMiner [373]. Point [106, 191, 322, 128]. Points [323, 306]. Poker [40]. Pole [225]. Policies [280, 102]. Polish [121, 123].

Poly [157]. Polynomial [47]. Polynomials [468]. Polyphonic [206]. Possibilistic [113]. possibility [69]. Post [164]. Post-Traumatic [164]. Potential [137, 38, 371]. Power [391, 471, 41, 165, 345]. Power-Law [165]. PPA [252]. Practical [508, 439, 404]. Precipitation [195]. precision [470]. Predicting [364]. Prediction [393, 350, 33, 148, 217, 154]. Predictions [155]. Predictive [217]. Prepositional [122]. Preprocessing [155]. Presence [450]. Present [64]. Preservation [52]. Preserving [66, 116, 309, 13]. Price [138, 408]. Prime [69]. Prime-Order [69]. PRINCE [77]. Principal [128]. Privacy [116, 309, 13, 295, 302]. Privacy-Preserving [309, 13]. Probabilistic [115, 48]. Probability [99]. Problem [318, 88, 327, 107, 437, 97, 333, 256, 6, 385, 205, 70, 322]. Problems [50, 202, 232, 325, 387, 331, 70]. Proceeding [259]. Proceedings [493, 494, 495, 496, 497, 487, 488, 489, 502, 492, 491, 498, 507, 504, 503, 505, 509, 506, 500, 490, 501, 499]. Process [171, 4, 156, 253]. Processing [224, 493, 494, 495, 496, 497, 102, 252, 503, 226]. Producing [424]. Production [150]. Profile [32, 455]. Profiles [2]. Profit [101]. Program [438, 442, 419]. programmable [475]. Programming [317]. Programs [73, 507, 435]. Progress [498]. Projecting [69]. Projections [62]. Promote [126, 210]. Proof [430, 432, 411, 475, 435, 439]. Proof-Assistant [439]. Proofs [418, 431, 507]. Propagation [36]. Properties [60, 15]. Property [430, 480, 478, 69]. Protection [296, 480, 485]. Protocol [286, 294, 281]. Protocol-Independent [281]. Protocols [237]. ProtoLeaks [281]. Proving [67, 440, 426]. Proximity [446, 357]. Proxy [476, 309]. Pseudo [243]. Pseudo-random [243]. Pseudoinversion [224]. PSO [389]. Psychological [264]. Public [462, 292, 470, 306]. Public-Key [470]. Publishing [5]. Pulse [222]. PVS [440]. PVT [15].

Qatar [493, 494, 495, 496, 497]. Quad [190]. Quadratic [468]. Quality [337, 349, 313, 373, 199]. Query [362, 402, 368]. Query-By-Object [402]. Query-Oriented [362]. Question [358]. Queueing [102]. Quickcheck [425].

Radial [394, 342]. Radio [36]. Rainbow [248]. Random [382, 100, 83, 475, 243]. Randomisations [116]. Randomised [97]. Randomized [74]. Ranging [208]. Rank [70]. Rare [111]. Rasterization [211, 212]. Rates [153]. Rating [434]. Raw [225]. RBAC [277]. RBF [148]. Re [309]. Re-encryption [309]. Reactions [119]. Reactive [391]. Real [95, 378, 206, 225, 352, 205, 344]. Real-Parameter [378]. Real-Time [206, 352, 205, 344]. Reasonable [242]. Reasoning [257]. Reassignment [376]. Rebound [246]. Recognition [157, 456, 46, 457, 200, 446, 462, 198, 254, 147, 352, 355, 214, 9, 336, 222]. Recognizing [460]. Recommendation [395, 5, 412, 374, 2, 4]. Reconstruction [62]. Records [195]. Rectangular [86, 55]. Reduced [244, 467]. Reduced-Round [244, 467]. Reducing [464, 345]. Reduction [475, 410, 192]. Reduction-Centric [475]. Reductions [70]. redundant [108]. Reference [355]. Refinement [109, 353]. Regional [56, 227]. Regression [387]. Regular [74]. Regulatory [321]. Reinforcement [214]. Related [464]. Relation [367, 27]. Relational [405]. Relations [116, 122]. Relationship [192, 123]. Relative [111]. Reliable [281]. RelSup [111]. Remanufacturing [379]. Remote [143]. Removal [169]. Renewable [392]. Reoptimization [88]. Replacement [116]. Replicating [271]. Represent [24]. Representation [163, 198, 336, 219]. Reproduced [165]. Reproduction [326]. Requirement [402]. Requirements [337]. Reranking [351]. Research [5]. Researcher [373]. Resilient [237]. Resistant [289]. Resolution [220]. Resources [205]. Response [451]. Responses [152]. Rest [136]. Retail [176]. Retention [116]. Retrieval [255, 504, 358, 375, 265, 214, 357]. Reusability [359]. Revelation [264]. Reversible [283]. Revised [510, 508]. Revisited [36, 83, 241, 72, 360]. Revocation [287]. Rewriting [56, 365]. Rigid [60]. Rim [503]. Risk [277, 155, 154]. Risk-Aware [277]. Robust [7, 216, 340, 306, 348]. Robustly [165]. ROI [223]. ROI-HOG [223]. Role [257]. Rooms [451, 463]. Root [435]. ROSETTA [239]. Rough [113, 204]. Round [244, 467]. Rounds [246]. Route [91]. Routing [87, 37, 390]. RPKM [113]. Rule [324, 111, 283, 112, 259]. Rule-Based [324]. Rules [108, 106, 112]. Rumor [96].

S [385]. Saddle [323]. SafeCode [295]. Safeguarding [295]. Safety [189, 501]. Sales [364]. Saliency [351, 227]. Salient [175]. Sammon [26]. Samples [271]. Sampling [201]. Satellite [384]. Satisfying [1]. Scalable [416, 399, 30]. Scale [7, 341, 190, 331]. Scaling [87, 165]. Scatters [229]. Scheduling [88, 101]. Schema [405]. Scheme [476, 309, 304]. Schemes [378, 71]. School [508]. SEAL [502]. Search [388, 327, 321, 202, 379, 371, 351, 135, 359, 32, 247, 361]. Searchable [307]. Second [507]. Secure [303, 73, 287, 71, 237, 294, 471, 234, 284, 311]. Security [274, 308, 67, 81, 475, 510, 275, 295, 500, 490, 466, 501]. Seeds [365, 365]. Segment [339]. Segment-Based [339]. Segmentation [58, 204, 168, 227, 214]. Selected [510]. Selection [93, 270, 171, 181, 413, 175]. Selective [397]. Self [207, 173, 292, 218, 162, 186, 145, 132]. Self-Adaptive [173]. Self-Organising [186]. Self-Organized [218, 292]. Self-Organizing [207, 162]. Sells [364]. Semantic [122, 365, 32]. Semantics [423, 417, 27]. SEMCCO [505]. Semi [457, 9]. Semi-supervised [457, 9]. Sensing [446]. Sensitivity [15]. Sensor [286, 384, 180]. Sensors [456, 445]. Sensory [152]. Sentiment [364]. Separating [45]. Separation [189]. Sequence [291]. Sequences [132]. Sequential [152, 6]. Series [170, 148, 217, 8]. Server [485]. Services [34, 33, 455]. Sessions [277]. Set [107, 266, 93, 365]. Sets [204, 177, 16, 61]. Setting [69]. Settling [369]. Several [213, 246]. Shall [429]. Shamir [235]. Shape [254, 223]. Shared [183]. Sharing [301]. Sheets [207]. Shift [205]. Short [35]. Short-Term [35]. Shortest [90, 95, 43]. Shortest-Elapsed-Time-First [90]. Side [483]. Signals [216, 268]. Signature [71, 473]. Signatures [66, 67]. Similarity [269, 349]. Simple [66]. Simulated [413, 168, 179, 502]. Simulation [332, 266, 98, 176]. Simulation-Based [98]. Simulations [334]. Simultaneous [213]. Singapore [503]. Single [239, 197, 246]. Single-SP [246]. SIRMs [160]. Size [66]. Sizing [394]. Skeletal [62]. Skill [256]. Sleep [266, 101, 269]. Slice [165]. Slicing [44]. Sliding [299]. Sliding-Mode [299]. Slow [101]. Small [50, 249, 176, 354, 165]. Small-World [176, 165]. Smart [303, 447, 460, 412, 458, 455, 463, 322, 302]. Smartphones [297]. Social [460, 96, 406, 3, 369, 300, 353, 126, 210]. Societies [504]. Soft [480]. Software [393, 273, 508, 439]. Solution [385, 376]. Solutions [315]. Solving [202, 256, 3, 205, 259]. Somatosensory [134]. Some [50, 70, 65]. Somewhat [311]. Sorting [318, 330, 376]. Sorting-Based [376]. Sound [458, 208]. Sound-Based [208]. Source [187, 189]. SP [246]. Space [50, 16, 131, 22]. Spaces [194]. Spanners [92]. Sparse [181, 198, 219]. Sparsest [107]. Sparsity [159]. Spatial [282, 347]. Spatio [35, 120, 24]. Spatio-temporal [35, 120, 24]. Speaker [133]. Specialising [457]. Specific [330, 183]. Specifications [216, 424]. Spectral [407]. Spectrometry [332]. Speech [147, 355]. Speed [87, 248]. Spiking [368]. SPL [37]. SPL-OBS [37]. Sponsored [371]. Spreading [96]. Square [435]. Squeezing [238]. SSH [311]. SSIM [346]. SSIM-Based [346]. Stages [269]. State [437, 174, 136]. States [134]. Static [67, 297]. Station [318]. Statistic [357]. Statistical [401, 319]. Statistics [261]. Status [137]. Steering [448]. Steganography [282]. Stiefel [139]. Stiffness [258]. Stimulus [136]. Stolen [295]. STPHD [180]. STPHD-Based [180]. Strategy [382, 397]. Stratified [18]. Stream [80]. Streaming [409]. Strengths [233]. Stress [164]. String [431, 333]. Stripes [354]. Structural [11, 349]. Structure [66, 374, 161, 13]. Structure-Preserving [66]. Structured [22]. Structuring [16, 51]. Study [327, 46, 255, 109, 371, 252, 256, 254, 479, 214, 212, 219]. Subjects [137]. Sudoku [259]. Sum [327]. Summaries [197]. Summarization [407]. Summarizing [14, 405]. Summer [508]. Super [220]. supervised [457, 9]. Supervision [186]. Support [447, 11, 117, 111, 182, 147, 455]. Supporting [308]. Surveillance [120]. SVMs [157]. SVR [148]. Swarm [386, 381, 392, 505]. Swing [225]. Switching [174]. Symmetric [72, 236]. Symposium [492, 501]. Synchronization [174, 161, 165]. Syndrome [262, 250]. Syndrome-Based [250]. Synthesis [382, 343]. System [193, 290, 200, 248, 206, 427, 394, 54, 151, 392, 221, 211, 212, 208, 201, 304, 342, 345, 194]. Systemic [274]. Systems [303, 492, 39, 480, 453, 20, 83, 203, 4, 15, 399, 449, 500, 141].

Tag [353]. tampering [105]. Tandem [332]. Target [93]. Targeted [111]. Task [183]. Tate [474]. Teaching [397]. Teaching-Learning [397]. Technique [397]. Techniques [196, 440]. Techno [392]. Techno-Economic [392]. Technologies [487, 488]. Technology [504]. Teddy [450]. Template [350]. Templates [223]. Temporal [262, 350, 35, 120, 24]. Tensor [11, 178]. Term [35]. Terminal [299]. Test [137, 254, 359]. Texas [40]. Text [121, 124, 366]. Texts [123]. Textual [407]. Texture [228]. Their [411, 450]. Thematic [363]. Theorem [440, 431]. Theoretic [270]. Theoretical [400, 126]. Theory [16, 490]. Thermal [149]. Thinning [52]. Third [505, 509]. Threat [288]. Three [471, 218, 112]. Threshold [484]. Thresholding [380]. Thresholds [93]. Tianjin [504]. tier [298]. Tight [81]. Tightly [45]. Tile [56]. Time [88, 90, 73, 337, 170, 148, 206, 217, 13, 352, 131, 205, 8, 344]. Timeline [479]. Tolerant [92, 390]. Tool [260]. Tools [508]. Top [108]. Top- [108]. Topology [98, 52, 174, 53, 165]. Total [88]. Trace [239]. Tracking [11, 383]. Tractable [93]. Traffic [310, 201]. TrafficS [201]. Training [457, 17]. Trajectory [13]. Transactional [25]. Transcoder [347]. Transfer [454]. Transform [235]. Transformation [372, 307]. Transformations [60]. Transient [131]. Transient-Time [131]. Transitions [361]. Transitive [433]. TRARM [111]. TRARM-RelSup [111]. Traumatic [164]. Tree [12]. Tree-Based [12]. Trees [170, 110, 111]. Triangular [52, 53]. Trigonometry [131]. Trimmed [144]. Triples [123]. Triviality [234]. Truncated [247]. Trust [278, 284, 279, 279]. Trust-orBAC [279]. Tsallis [380]. Tumor [127]. Turning [35]. Tutorial [508]. TV [412]. Tweets [369]. Twitter [370, 369, 368]. Two [238, 62, 333, 1, 72]. Two-Level [1]. Two-Party [72]. Two-Phase [333]. TX [489]. Type [291, 15]. Type- [15].

Undergoing [137]. Understanding [119, 451, 83, 132]. Unified [234]. Unit [222]. Unit-Linking [222]. Units [360]. UOWHFs [74]. Update [286]. USA [489]. Use [150]. User [447, 370, 460, 2, 295, 449, 32, 30, 463]. Users [369]. Using [388, 34, 157, 386, 87, 305, 393, 391, 337, 456, 350, 457, 200, 480, 445, 68, 442, 7, 216, 195, 207, 44, 138, 233, 408, 413, 111, 198, 2, 190, 118, 160, 410, 28, 197, 394, 3, 389, 282, 339, 122, 268, 150, 331, 384, 312, 183, 306, 192, 392, 363, 300, 132, 208, 228, 130, 159, 222, 199, 342]. Utility [12].

V [497]. V2GPriv [302]. V2V [287]. Validation [156]. Valley [323]. Valley-Adaptive [323]. Vandalism [372]. Variable [469]. Variance [94]. Variational [177]. Vector [11, 182, 147]. Vectorial [243]. Vehicle [302, 222]. Vehicle-to-Grid [302]. Vein [149]. verbal [450]. Verification [290, 417, 118, 508, 439, 441, 428, 421]. Verified [420]. Verifier [442]. Version [392]. Vertex [97]. Veto [41]. via [92, 171, 261, 3, 353, 223]. Victims [481]. Video [350, 346]. Videoconference [342]. Videoconferencing [347]. Videoconferencing-Oriented [347]. Videos [336, 221]. Vietnam [502, 487, 488]. Vietnamese [9]. View [413, 106, 343]. Viewing [260]. Virtual [404, 428]. Visibility [269]. Vision [254, 191]. Visual [133, 163, 225, 136, 135, 162, 258, 158, 456]. Visualisation [163]. Visualization [310, 26, 479, 145]. Visuomotor [258]. Vote [20]. Voxel [270]. VulnerableMe [274].

Wall [464]. Wandering [464]. Warehouse [23]. Watermarking [283]. Watershed [58]. Watershed-Based [58]. Way [74]. Weaker [246]. Weaknesses [274]. Web [34, 35, 482, 33, 485, 367, 479, 32, 455, 30, 361]. Weighted [12]. Weighting [20]. Wheat [150]. Wheel [448]. Where [369]. White [354]. Wide [342]. Wikipedia [372, 365]. Wireless [98, 285]. WISA [510]. Wise [59]. Wizard [459]. Word [157, 247]. Word-Oriented [247]. Words [336]. Work [50]. Workshop [489, 510]. Workspace [389]. World [176, 165]. Wormhole [285].

XCS [328].

Young [450].

Zero [79]. Zone [141]. ZUC [80].

References

Kozierkiewicz-Hetmanska:2012:COL

[1] Adrianna Kozierkiewicz-Hetmańska. Comparison of one-level and twolevel consensuses satisfying the 2-optimality criterion. Lecture Notes in Computer Science, 7653:1–10, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

     chapter/10.1007/978-3-642-34630-9_1/.

Maleszka:2012:HMC

[2] Marcin Maleszka, Bernadetta Mianowska, and Ngoc-Thanh Nguyen. A heuristic method for collaborative recommendation using hierarchical user profiles. Lecture Notes in Computer Science, 7653:11–20, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

     link.springer.com/chapter/10.1007/978-3-642-34630-9_2/.

Nguyen:2012:SCC

[3] Quoc Uy Nguyen, Trong Hai Duong, and Sanggil Kang. Solving conflict on collaborative knowledge via social networking using consensus choice. Lecture Notes in Computer Science, 7653:21–30, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

     link.springer.com/chapter/10.1007/978-3-642-34630-9_3/.

Pham:2012:IME

[4] Xuan Hau Pham, Jason J. Jung, and Ngoc-Thanh Nguyen. Integrating multiple experts for correction process in interactive recommendation systems. Lecture Notes in Computer Science, 7653:31–40, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34630-9_4/.

Huynh:2012:MCK

[5] Tin Huynh and Kiem Hoang. Modeling collaborative knowledge of publishing activities for research recommendation. Lecture Notes in Computer Science, 7653:41–50, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34630-9_5/.

Nguyen:2012:NAP

[6] Thanh-Trung Nguyen and Phi-Khu Nguyen. A new approach for problem of sequential pattern mining. Lecture Notes in Computer Science, 7653:51–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34630-9_6/.

Hoang:2012:RHD

[7] Van-Dung Hoang, My-Ha Le, and Kang-Hyun Jo. Robust human detection using multiple scale of cell based histogram of oriented gradients and AdaBoost learning. Lecture Notes in Computer Science, 7653: 61–71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34630-9_7/.

Son:2012:DTS

[8] Nguyen Thanh Son and Duong Tuan Anh. Discovering time series motifs based on multidimensional index and early abandoning. Lecture Notes in Computer Science, 7653:72–82, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34630-9_8/.

Vo:2012:HAP

[9] Duc-Thuan Vo and Cheol-Young Ock. A hybrid approach of pattern extraction and semi-supervised learning for Vietnamese named entity recognition. Lecture Notes in Computer Science, 7653:83–93, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34630-9_9/.

Pawlikowski:2012:IEG

[10] Roman Pawlikowski, Krzysztof Ociepa, Urszula Markowska-Kaczmar, and Pawel B. Myszkowski. Information extraction from geographical overview maps. Lecture Notes in Computer Science, 7653:94–103, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34630-9_10/.

Cyganek:2012:PBO

[11] Bogusław Cyganek and Michał Woźniak. Pixel-based object detection and tracking with ensemble of support vector machines and extended structural tensor. Lecture Notes in Computer Science, 7653:104–113, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34630-9_11/.

Vo:2012:TBA

[12] Bay Vo, Bac Le, and Jason J. Jung. A tree-based approach for mining frequent weighted utility itemsets. Lecture Notes in Computer Science, 7653:114–123, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-34630-9_12/.

Phan:2012:NTP

[13] Trong Nhan Phan and Tran Khanh Dang. A novel trajectory privacypreserving future time index structure in moving object databases. Lecture Notes in Computer Science, 7653:124–134, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34630-9_13/.

Skorupa:2012:SKB

[14] Grzegorz Skorupa and Radosław Katarzyniak. Summarizing knowledge base with modal conditionals. Lecture Notes in Computer Science, 7653: 135–144, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34630-9_14/.

Selamat:2012:MPP

[15] Ali Selamat, S. O. Olatunji, and Abdul Azeez Abdul Raheem. Modeling PVT properties of crude oil systems based on type-2 fuzzy logic approach and sensitivity based linear learning method. Lecture Notes in Computer Science, 7653:145–155, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34630-9_15/.

Jastrzebska:2012:SSN

[16] Agnieszka Jastrzebska and Wladyslaw Homenda. On structuring of the space of needs in the framework of fuzzy sets theory. Lecture Notes in Computer Science, 7653:156–165, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34630-9_16/.

Wilk:2012:CFC

[17] Tomasz Wilk and Michał Woźniak. Comparison of fuzzy combiner training methods. Lecture Notes in Computer Science, 7653:166–173, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34630-9_17/.

Tran:2012:AMM

[18] Trong Hieu Tran and Quoc Bao Vo. An axiomatic model for merging stratified belief bases by negotiation. Lecture Notes in Computer Science, 7653:174–184, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-34630-9_18/.

Pedrycz:2012:FCM

[19] Witold Pedrycz and Wladyslaw Homenda. From fuzzy cognitive maps to granular cognitive maps. Lecture Notes in Computer Science, 7653: 185–193, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34630-9_19/.

Hardas:2012:BVW

[20] Manas S. Hardas and Lisa Purvis. Bayesian vote weighting in crowdsourcing systems. Lecture Notes in Computer Science, 7653:194–203, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34630-9_20/.

Anonymous:2012:FMa

[21] Anonymous. Front matter. Lecture Notes in Computer Science, 7653: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-34630-9/1.

Sitarek:2012:GOS

[22] Tomasz Sitarek and Wladyslaw Homenda. Generic operations in the structured space of the music. Lecture Notes in Computer Science, 7654:1–10, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34707-8_1/.

Hoang:2012:CCP

[23] Duong Thi Anh Hoang, Ngoc Sy Ngo, and Binh Thanh Nguyen. Collective cubing platform towards definition and analysis of warehouse cubes. Lecture Notes in Computer Science, 7654:11–20, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34707-8_2/.

Tran:2012:ACC

[24] Phuoc Vinh Tran. To approach cylindrical coordinates to represent multivariable spatio-temporal data. Lecture Notes in Computer Science, 7654: 21–28, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34707-8_3/.

Herawan:2012:EME

[25] Tutut Herawan, A. Noraziah, Zailani Abdullah, Mustafa Mat Deris, and Jemal H. Abawajy. EFP-M2: Efficient model for mining frequent patterns in transactional database. Lecture Notes in Computer Science, 7654:29–38, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34707-8_4/.

Kwasnicka:2012:ISM

[26] Halina Kwasnicka and Pawel Siemionko. Improved Sammon mapping method for visualization of multidimensional data. Lecture Notes in Computer Science, 7654:39–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34707-8_5/.

Pietranik:2012:ORA

[27] Marcin Mirosław Pietranik and Ngoc Thanh Nguyen. Ontology relation alignment based on attribute semantics. Lecture Notes in Computer Science, 7654:49–58, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-34707-8_6/.

Moon:2012:DDU

[28] Young Chan Moon, Ho Min Jung, Chuck Yoo, and Young Woong Ko. Data deduplication using dynamic chunking algorithm. Lecture Notes in Computer Science, 7654:59–68, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34707-8_7/.

Dang:2012:AMF

[29] Huynh Tu Dang, Ha Manh Tran, Phach Ngoc Vu, and An Truong Nguyen. Applying MapReduce framework to peer-to-peer computing applications. Lecture Notes in Computer Science, 7654:69–78, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34707-8_8/.

Wecel:2012:SAW

[30] Krzysztof Wȩcel, Tomasz Kaczmarek, and Agata Filipowska. Scalable adaptation of Web applications to users’ behavior. Lecture Notes in Computer Science, 7654:79–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34707-8_9/.

Andres:2012:OOC

[31] César Andrés, Rui Abreu, and Alberto Núñez. OCE: An online colaborative Editor. Lecture Notes in Computer Science, 7654:89–98, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34707-8_10/.

Uddin:2012:CSU

[32] Mohammed Nazim Uddin, Trong Hai Duong, Visal Sean, and Geun-Sik Jo. Construction of semantic user profile for personalized Web search. Lecture Notes in Computer Science, 7654:99–108, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34707-8_11/.

Grzech:2012:LPD

[33] Adam Grzech, Krzysztof Juszczyszyn, Paweł Stelmach, and Łukasz Falas. Link prediction in dynamic networks of services emerging during deployment and execution of Web services. Lecture Notes in Computer Science, 7654:109–120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-34707-8_12/.

Al-Zyoud:2012:TMC

[34] Mahran Al-Zyoud, Imad Salah, and Nadim Obeid. Towards a model of context awareness using Web services. Lecture Notes in Computer Science, 7654:121–131, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-34707-8_13/.

Borzemski:2012:STS

[35] Leszek Borzemski, Michal Danielak, and Anna Kaminska-Chuchmala. Short-term spatio-temporal forecasts of Web performance by means of turning bands method. Lecture Notes in Computer Science, 7654:132– 141, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34707-8_14/.

Blaskiewicz:2012:EPA

[36] Przemysław Błaśkiewicz, Mirosław Kutyłowski, Wojciech Wodo, and Kamil Wolny. Extreme propagation in an ad-hoc radio network- revisited. Lecture Notes in Computer Science, 7654:142–151, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34707-8_15/.

Chuong:2012:MPA

[37] Dang Thanh Chuong, Vu Duy Loi, and Vo Viet Minh Nhat. A model for the performance analysis of SPL-OBS core nodes with deflection routing. Lecture Notes in Computer Science, 7654:152–161, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34707-8_16/.

Encheva:2012:OPC

[38] Sylvia Encheva. Ordering of potential collaboration options. Lecture Notes in Computer Science, 7654:162–171, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34707-8_17/.

Dong:2012:IDD

[39] Ching-Shen Dong and Ananth Srinivasan. Interface design for decision systems. Lecture Notes in Computer Science, 7654:172–181, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34707-8_18/.

Fedczyszyn:2012:OMT

[40] Grzegorz Fedczyszyn, Leszek Koszalka, and Iwona Pozniak-Koszalka. Opponent modeling in Texas hold’em poker. Lecture Notes in Computer Science, 7654:182–191, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34707-8_19/.

Mercik:2012:API

[41] Jacek Mercik. On axiomatization of power index of veto. Lecture Notes in Computer Science, 7654:192–200, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34707-8_20/.

Anonymous:2012:FMb

[42] Anonymous. Front matter. Lecture Notes in Computer Science, 7654: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-34707-8/1.

Dutt:2012:FSI

[43] Mousumi Dutt, Arindam Biswas, Partha Bhowmick, and Bhargab B. Bhattacharya. On finding shortest isothetic path inside a digital object. Lecture Notes in Computer Science, 7655:1–15, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.com/chapter/10.1007/978-3-642-34732-0_1/.

Karmakar:2012:FSO

[44] Nilanjana Karmakar, Arindam Biswas, and Partha Bhowmick. Fast slicing of orthogonal covers using DCEL. Lecture Notes in Computer Science, 7655:16–30, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34732-0_2/.

Veelaert:2012:FCA

[45] Peter Veelaert. Fast combinatorial algorithm for tightly separating hyperplanes. Lecture Notes in Computer Science, 7655:31–44, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34732-0_3/.

Chen:2012:DCA

[46] Li Chen and Soma Biswas. Digital curvatures applied to 3D object analysis and recognition: a case study. Lecture Notes in Computer Science, 7655:45–58, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34732-0_4/.

Sekiya:2012:DPC

[47] Fumiki Sekiya and Akihiro Sugimoto. Discrete polynomial curve fitting to noisy data. Lecture Notes in Computer Science, 7655:59–74, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34732-0_5/.

Herrera-Navarro:2012:PMC

[48] Ana Marcela Herrera-Navarro, Hugo Jiménez-Hernández, and Iván Ramón Terol-Villalobos. A probabilistic measure of circularity. Lecture Notes in Computer Science, 7655:75–89, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34732-0_6/.

Asano:2012:NFC

[49] Tetsuo Asano and Sergey Bereg. A new framework for connected components labeling of binary images. Lecture Notes in Computer Science, 7655:90–102, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34732-0_7/.

Asano:2012:SWS

[50] Tetsuo Asano, Sergey Bereg, and Lilian Buzer. Small work space algorithms for some basic problems on binary images. Lecture Notes in Computer Science, 7655:103–114, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34732-0_8/.

Slapal:2012:ASD

[51] Josef Šlapal. Adjacencies for structuring the digital plane. Lecture Notes in Computer Science, 7655:115–127, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34732-0_9/.

Kardos:2012:TPT

[52] Péter Kardos and Kálmán Palágyi. On topology preservation for triangular thinning algorithms. Lecture Notes in Computer Science, 7655: 128–142, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34732-0_10/.

Nagy:2012:CTT

[53] Benedek Nagy. Cellular topology on the triangular grid. Lecture Notes in Computer Science, 7655:143–153, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34732-0_11/.

Subramanian:2012:PSM

[54] K. G. Subramanian, Ibrahim Venkat, and Petra Wiederhold. A P system model for contextual array languages. Lecture Notes in Computer Science, 7655:154–165, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-34732-0_12/.

Lalitha:2012:RAP

[55] D. Lalitha, K. Rangarajan, and Durairaj Gnanaraj Thomas. Rectangular arrays and Petri nets. Lecture Notes in Computer Science, 7655: 166–180, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34732-0_13/.

Kamaraj:2012:RHT

[56] Thangasamy Kamaraj and Durairaj Gnanaraj Thomas. Regional hexagonal tile rewriting grammars. Lecture Notes in Computer Science, 7655: 181–195, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34732-0_14/.

Kamaraj:2012:PCA

[57] Thangasamy Kamaraj, Durairaj Gnanaraj Thomas, H. Geetha, and T. Kalyani. Partial commutation on array languages. Lecture Notes in Computer Science, 7655:196–208, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34732-0_15/.

Attig:2012:ILM

[58] Anja Attig and Petra Perner. Incremental learning of the model for watershed-based image segmentation. Lecture Notes in Computer Science, 7655:209–222, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34732-0_16/.

Gonzalez:2012:FLW

[59] Damien Gonzalez, Rémy Malgouyres, Henri-Alex Esbelin, and Chafik Samir. Fast level-wise convolution. Lecture Notes in Computer Science, 7655:223–233, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-34732-0_17/.

Ngo:2012:CPD

[60] Phuc Ngo, Yukiko Kenmochi, Nicolas Passat, and Hugues Talbot. Combinatorial properties of 2D discrete rigid transformations under pixelinvariance constraints. Lecture Notes in Computer Science, 7655:234– 248, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34732-0_18/.

Potluri:2012:NMA

[61] Anupama Potluri and Chakravarthy Bhagvati. Novel morphological algorithms for dominating sets on graphs with applications to image analysis. Lecture Notes in Computer Science, 7655:249–262, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34732-0_19/.

Hantos:2012:BIR

[62] Norbert Hantos, Péter Balázs, and Kálmán Palágyi. Binary image reconstruction from two projections and skeletal information. Lecture Notes in Computer Science, 7655:263–273, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34732-0_20/.

Anonymous:2012:FMc

[63] Anonymous. Front matter. Lecture Notes in Computer Science, 7655: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-34732-0/1.

Boneh:2012:PBC

[64] Dan Boneh. Pairing-based cryptography: Past, present, and future. Lecture Notes in Computer Science, 7658:1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/accesspage/chapter/10.1007/978-3-642-34961-4_1.

Zong:2012:SMM

[65] Chuanming Zong. Some mathematical mysteries in lattices. Lecture Notes in Computer Science, 7658:2–3, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

accesspage/chapter/10.1007/978-3-642-34961-4_2.

Abe:2012:CSS

[66] Masayuki Abe, Melissa Chase, Bernardo David, Markulf Kohlweiss, and Ryo Nishimaki. Constant-size structure-preserving signatures: Generic constructions and simple assumptions. Lecture Notes in Computer Science, 7658:4–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-34961-4_3/.

Gerbush:2012:DFS

[67] Michael Gerbush, Allison Lewko, Adam O’Neill, and Brent Waters. Dual form signatures: An approach for proving security from static assumptions. Lecture Notes in Computer Science, 7658:25–42, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34961-4_4/.

Hayashi:2012:BPB

[68] Takuya Hayashi, Takeshi Shimoyama, Naoyuki Shinohara, and Tsuyoshi Takagi. Breaking pairing-based cryptosystems using η_{T} pairing over GF(3^{97}). Lecture Notes in Computer Science, 7658:43–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34961-4_5/.

Seo:2012:IPP

[69] Jae Hong Seo. On the (im)possibility of projecting property in prime-order setting. Lecture Notes in Computer Science, 7658:61–79, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34961-4_6/.

Villar:2012:ORS

[70] Jorge Luis Villar. Optimal reductions of some decisional problems to the rank problem. Lecture Notes in Computer Science, 7658:80–97, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34961-4_7/.

Faust:2012:SSS

[71] Sebastian Faust, Carmit Hazay, Jesper Buus Nielsen, and Peter Sebastian Nordholt. Signature schemes secure against hard-to-invert leakage. Lecture Notes in Computer Science, 7658:98–115, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34961-4_8/.

Lindell:2012:CST

[72] Yehuda Lindell, Eran Omri, and Hila Zarosim. Completeness for symmetric two-party functionalities- revisited. Lecture Notes in Computer Science, 7658:116–133, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34961-4_9/.

Bellare:2012:ASG

[73] Mihir Bellare, Viet Tung Hoang, and Phillip Rogaway. Adaptively secure garbling with applications to one-time programs and secure outsourcing. Lecture Notes in Computer Science, 7658:134–153, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34961-4_10/.

Ames:2012:GRI

[74] Scott Ames, Rosario Gennaro, and Muthuramakrishnan Venkitasubramaniam. The generalized randomized iterate and its application to new efficient constructions of UOWHFs from regular one-way functions. Lecture Notes in Computer Science, 7658:154–171, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34961-4_11/.

Liu:2012:PAI

[75] Meicheng Liu, Yin Zhang, and Dongdai Lin. Perfect algebraic immune functions. Lecture Notes in Computer Science, 7658:172–189, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34961-4_12/.

Mendel:2012:DAL

[76] Florian Mendel, Vincent Rijmen, Deniz Toz, and Kerem Varıcı. Differential analysis of the LED block cipher. Lecture Notes in Computer Science, 7658:190–207, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34961-4_13/.

Borghoff:2012:PLL

[77] Julia Borghoff, Anne Canteaut, Tim Güneysu, Elif Bilge Kavun, and Miroslav Knezevic. PRINCE — a low-latency block cipher for pervasive computing applications. Lecture Notes in Computer Science, 7658: 208–225, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34961-4_14/.

Leurent:2012:ADA

[78] Gaëtan Leurent. Analysis of differential attacks in ARX constructions. Lecture Notes in Computer Science, 7658:226–243, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34961-4_15/.

Bogdanov:2012:IML

[79] Andrey Bogdanov, Gregor Leander, Kaisa Nyberg, and Meiqin Wang. Integral and multidimensional linear distinguishers with correlation zero. Lecture Notes in Computer Science, 7658:244–261, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.com/chapter/10.1007/978-3-642-34961-4_16/.

Wu:2012:DAA

[80] Hongjun Wu, Tao Huang, Phuong Ha Nguyen, Huaxiong Wang, and San Ling. Differential attacks against stream cipher ZUC. Lecture Notes in Computer Science, 7658:262–277, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34961-4_17/.

Lampe:2012:ATS

[81] Rodolphe Lampe, Jacques Patarin, and Yannick Seurin. An asymptotically tight security analysis of the iterated even-Mansour cipher. Lecture Notes in Computer Science, 7658:278–295, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34961-4_18/.

Zhang:2012:EMB

[82] Liting Zhang, Wenling Wu, Han Sui, and Peng Wang. 3kf9: Enhancing 3GPP-MAC beyond the birthday bound. Lecture Notes in Computer Science, 7658:296–312, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34961-4_19/.

Jetchev:2012:UAR

[83] Dimitar Jetchev, Onur Özen, and Martijn Stam. Understanding adaptivity: Random systems revisited. Lecture Notes in Computer Science, 7658: 313–330, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34961-4_20/.

Anonymous:2012:FMd

[84] Anonymous. Front matter. Lecture Notes in Computer Science, 7658: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-34961-4/1.

Amir:2012:DAP

[85] Amihood Amir, Alberto Apostolico, and Estrella Eisenberg. Detecting approximate periodic patterns. Lecture Notes in Computer Science, 7659:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34862-4_1/.

Ballard:2012:GEA

[86] Grey Ballard, James Demmel, and Olga Holtz. Graph expansion analysis for communication costs of fast rectangular matrix multiplication. Lecture Notes in Computer Science, 7659:13–36, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34862-4_2/.

Bansal:2012:MRE

[87] Nikhil Bansal, Anupam Gupta, and Ravishankar Krishnaswamy. Multicast routing for energy minimization using speed scaling. Lecture Notes in Computer Science, 7659:37–51, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34862-4_3/.

Baram:2012:RMT

[88] Guy Baram and Tami Tamir. Reoptimization of the minimum total flow-time scheduling problem. Lecture Notes in Computer Science, 7659: 52–66, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34862-4_4/.

Barcelo:2012:EEC

[89] Neal Barcelo, Miao Zhou, Daniel Cole, and Michael Nugent. Energy efficient caching for phase-change memory. Lecture Notes in Computer Science, 7659:67–81, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34862-4_5/.

Barcelo:2012:SET

[90] Neal Barcelo, Sungjin Im, Benjamin Moseley, and Kirk Pruhs. Shortestelapsed-time-first on a multiprocessor. Lecture Notes in Computer Science, 7659:82–92, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34862-4_6/.

Batz:2012:ERC

[91] Gernot Veit Batz, Robert Geisberger, and Dennis Luxen. Efficient route compression for hybrid route planning. Lecture Notes in Computer Science, 7659:93–107, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34862-4_7/.

Chechik:2012:MSF

[92] Shiri Chechik, Quentin Godfroy, and David Peleg. Multipath spanners via fault-tolerant spanners. Lecture Notes in Computer Science, 7659: 108–119, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34862-4_8/.

Chopin:2012:CTC

[93] Morgan Chopin, André Nichterlein, and Rolf Niedermeier. Constant thresholds can make target set selection tractable. Lecture Notes in Computer Science, 7659:120–133, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34862-4_9/.

Cunial:2012:FVC

[94] Fabio Cunial. Faster variance computation for patterns with gaps. Lecture Notes in Computer Science, 7659:134–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34862-4_10/.

DAngelo:2012:ECD

[95] Gianlorenzo D’Angelo and Mattia D’Emidio. Enhancing the computation of distributed shortest paths on real dynamic networks. Lecture Notes in Computer Science, 7659:148–158, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34862-4_11/.

Doerr:2012:EAR

[96] Benjamin Doerr, Mahmoud Fouz, and Tobias Friedrich. Experimental analysis of rumor spreading in social networks. Lecture Notes in Computer Science, 7659:159–173, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34862-4_12/.

ElOuali:2012:RAA

[97] Mourad El Ouali, Helena Fohlin, and Anand Srivastav. A randomised approximation algorithm for the partial vertex cover problem in hypergraphs. Lecture Notes in Computer Science, 7659:174–187, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34862-4_13/.

Fuchs:2012:SBA

[98] Fabian Fuchs, Markus Völker, and Dorothea Wagner. Simulation-based analysis of topology control algorithms for wireless ad hoc networks. Lecture Notes in Computer Science, 7659:188–202, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-34862-4_14/.

Goodrich:2012:COD

[99] Michael T. Goodrich and Daniel S. Hirschberg. Cache-oblivious dictionaries and multimaps with negligible failure probability. Lecture Notes in Computer Science, 7659:203–218, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34862-4_15/.

Gorke:2012:EGC

[100] Robert Görke, Roland Kluge, and Andrea Schumm. An efficient generator for clustered dynamic random networks. Lecture Notes in Computer Science, 7659:219–233, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34862-4_16/.

Kling:2012:SSP

[101] Peter Kling and Andreas Cord-Landwehr. Slow down and sleep for profit in online deadline scheduling. Lecture Notes in Computer Science, 7659: 234–247, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34862-4_17/.

Kogan:2012:FQP

[102] Kirill Kogan and Alejandro López-Ortiz. FIFO queueing policies for packets with heterogeneous processing. Lecture Notes in Computer Science, 7659:248–260, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-34862-4_18/.

Anonymous:2012:BMa

[103] Anonymous. Back matter. Lecture Notes in Computer Science, 7659: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-34862-4/1.

Anonymous:2012:FMe

[104] Anonymous. Front matter. Lecture Notes in Computer Science, 7659: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-34862-4/1.

Moio:2012:ATA

[105] Andrea Moio, Attilio Giordana, and Dino Mendola. An anti-tampering algorithm based on an artificial intelligence approach. Lecture Notes in Computer Science, 7661:1–10, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34624-8_1/.

Rauch:2012:DKD

[106] Jan Rauch. Domain knowledge and data mining with association rules — a logical point of view. Lecture Notes in Computer Science, 7661: 11–20, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34624-8_2/.

Betlinski:2012:PFS

[107] Paweł Betliński and Dominik Ślȩzak. The problem of finding the sparsest Bayesian network for an input data set is NP-Hard. Lecture Notes in Computer Science, 7661:21–30, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34624-8_3/.

Fournier-Viger:2012:MTN

[108] Philippe Fournier-Viger and Vincent S. Tseng. Mining top-K nonredundant association rules. Lecture Notes in Computer Science, 7661: 31–40, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34624-8_4/.

Guid:2012:AKR

[109] Matej Guid, Martin Možina, Vida Groznik, Dejan Georgiev, and Aleksander Sadikov. ABML knowledge refinement loop: a case study. Lecture Notes in Computer Science, 7661:41–50, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34624-8_5/.

Lavergne:2012:MMI

[110] Jennifer Lavergne, Ryan Benton, and Vijay V. Raghavan. Min-max itemset trees for dense and categorical datasets. Lecture Notes in Computer Science, 7661:51–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34624-8_6/.

Lavergne:2012:TRT

[111] Jennifer Lavergne, Ryan Benton, and Vijay V. Raghavan. TRARM-RelSup: Targeted rare association rule mining using itemset trees and the relative support measure. Lecture Notes in Computer Science, 7661: 61–70, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34624-8_7/.

Tsumoto:2012:IRI

[112] Shusaku Tsumoto and Shoji Hirano. Incremental rules induction method based on three rule layers. Lecture Notes in Computer Science, 7661: 71–80, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34624-8_8/.

Ammar:2012:RRP

[113] Asma Ammar, Zied Elouedi, and Pawan Lingras. RPKM: The rough possibilistic K-modes. Lecture Notes in Computer Science, 7661:81– 86, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34624-8_9/.

Kohout:2012:EEO

[114] Jan Kohout and Roman Neruda. Exploration and exploitation operators for genetic graph clustering algorithm. Lecture Notes in Computer Science, 7661:87–92, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34624-8_10/.

Clark:2012:LPA

[115] Patrick G. Clark, Jerzy W. Grzymala-Busse, and Martin Kuehnhausen. Local probabilistic approximations for incomplete data. Lecture Notes in Computer Science, 7661:93–98, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34624-8_11/.

Andruszkiewicz:2012:RBR

[116] Piotr Andruszkiewicz. On the relations between retention replacement, additive perturbation, and randomisations for nominal attributes in privacy preserving data mining. Lecture Notes in Computer Science, 7661: 99–104, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34624-8_12/.

Ghosh:2012:DSE

[117] Sujata Ghosh, Thiri Haymar Kyaw, and Rineke Verbrugge. Decision support for extensive form negotiation games. Lecture Notes in Computer Science, 7661:105–114, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34624-8_13/.

Meng:2012:IAF

[118] Yuxin Meng, Wenjuan Li, and Lam for Kwok. Intelligent alarm filter using knowledge-based alert verification in network intrusion detection. Lecture Notes in Computer Science, 7661:115–124, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34624-8_14/.

Gao:2012:CBM

[119] Chao Gao and Jiming Liu. Clustering-based media analysis for understanding human emotional reactions in an extreme event. Lecture Notes in Computer Science, 7661:125–135, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34624-8_15/.

Pestana:2012:EDA

[120] Gabriel Pestana, Joachim Metter, Sebastian Heuchler, and Pedro Reis. An event-driven architecture for spatio-temporal surveillance of business activities. Lecture Notes in Computer Science, 7661:136–142, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34624-8_16/.

Brocki:2012:MMT

[121] Łukasz Brocki, Krzysztof Marasek, and Danijel Koržinek. Multiple model text normalization for the Polish language. Lecture Notes in Computer Science, 7661:143–148, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34624-8_17/.

Punuru:2012:DSR

[122] Janardhana Punuru and Jianhua Chen. Discovering semantic relations using prepositional phrases. Lecture Notes in Computer Science, 7661: 149–154, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34624-8_18/.

Wroblewska:2012:DDB

[123] Alina Wróblewska and Marcin Sydow. DEBORA: Dependency-based method for extracting entity-relationship triples from open-domain texts in Polish. Lecture Notes in Computer Science, 7661:155–161, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34624-8_19/.

Protaziuk:2012:LOL

[124] Grzegorz Protaziuk, Anna Wróblewska, Robert Bembenik, and Henryk Rybiński. Lexical ontology layer — a bridge between text and concepts. Lecture Notes in Computer Science, 7661:162–171, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34624-8_20/.

Anonymous:2012:FMf

[125] Anonymous. Front matter. Lecture Notes in Computer Science, 7661: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-34624-8/1.

Xiao:2012:DSNa

[126] Lu Xiao and Ming Fan. Does social network always promote entrepreneurial intentions? Part I: Theoretical model. Lecture Notes in Computer Science, 7663:1–7, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34475-6_1/.

Feroz:2012:AAH

[127] Samreen Feroz, Amatal Habib, Maryam Siddiqua, Sobia Saleem, and Nisar Ahmed Shar. Association of anti-histamine drugs with brain tumor. Lecture Notes in Computer Science, 7663:8–15, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.com/chapter/10.1007/978-3-642-34475-6_2/.

Zhu:2012:EPP

[128] Qiuyu Zhu. Estimating principal point and nonlinear parameters of camera from a planar calibration image. Lecture Notes in Computer Science, 7663:16–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34475-6_3/.

Arora:2012:DDI

[129] Vipul Arora and Laxmidhar Behera. Design of distribution independent noise filters with online PDF estimation. Lecture Notes in Computer Science, 7663:25–32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-34475-6_4/.

Wu:2012:DEC

[130] Xia Wu, Juan Li, and Li Yao. Determining effective connectivity from FMRI data using a Gaussian dynamic Bayesian network. Lecture Notes in Computer Science, 7663:33–39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34475-6_5/.

Radwan:2012:TTF

[131] A. G. Radwan and Ahmed S. Elwakil. Transient-time fractional-space trigonometry and application. Lecture Notes in Computer Science, 7663: 40–47, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34475-6_6/.

Wickramasinghe:2012:UIP

[132] Manjusri Wickramasinghe, Jayantha Rajapakse, and Damminda Alahakoon. Understanding individual play sequences using growing self organizing maps. Lecture Notes in Computer Science, 7663:48–55, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34475-6_7/.

Almaadeed:2012:AVF

[133] Noor Almaadeed, Amar Aggoun, and Abbes Amira. Audio-visual feature fusion for speaker identification. Lecture Notes in Computer Science, 7663:56–67, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34475-6_8/.

Kang:2012:DCS

[134] Xiaoxu Kang, Marc Schieber, and Nitish V. Thakor. Decoding cognitive states from neural activities of somatosensory cortex. Lecture Notes in Computer Science, 7663:68–75, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34475-6_9/.

Neokleous:2012:CMD

[135] Kleanthis C. Neokleous, Marios N. Avraamides, Costas K. Neocleous, and Christos N. Schizas. Cognitive modeling of dilution effects in visual search. Lecture Notes in Computer Science, 7663:76–83, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34475-6_10/.

Mumtaz:2012:CAE

[136] Wajid Mumtaz, Likun Xia, Aamir Saeed Malik, and Mohd Azhar Mohd Yasin. Complexity analysis of EEG data during rest state and visual stimulus. Lecture Notes in Computer Science, 7663:84–91, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34475-6_11/.

Al-Ani:2012:AAS

[137] Ahmed Al-Ani, Bram Van Dun, Harvey Dillon, and Alaleh Rabie. Analysis of alertness status of subjects undergoing the cortical auditory evoked potential hearing test. Lecture Notes in Computer Science, 7663:92– 99, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34475-6_12/.

Kaur:2012:PFU

[138] Preetinder Kaur, Madhu Goyal, and Jie Lu. Price forecasting using dynamic assessment of market conditions and Agent’s bidding behavior. Lecture Notes in Computer Science, 7663:100–108, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34475-6_13/.

Kanamori:2012:NCO

[139] Takafumi Kanamori and Akiko Takeda. Non-convex optimization on Stiefel manifold and applications to machine learning. Lecture Notes in Computer Science, 7663:109–116, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34475-6_14/.

Dinu:2012:LPD

[140] Liviu Petrisor Dinu, Radu-Tudor Ionescu, and Marius Popescu. Local patch dissimilarity for images. Lecture Notes in Computer Science, 7663: 117–126, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34475-6_15/.

Zhang:2012:NOC

[141] Dehua Zhang, Derong Liu, and Qinglai Wei. Nearly optimal control for nonlinear systems with dead-zone control input based on the iterative ADP approach. Lecture Notes in Computer Science, 7663:127– 134, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34475-6_16/.

Benkhider:2012:MAK

[142] Sadjia Benkhider, Oualid Dahmri, and Habiba Drias. A memetic approach for the knowledge extraction. Lecture Notes in Computer Science, 7663: 135–141, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34475-6_17/.

Liu:2012:DNC

[143] Xiaohua Liu, Haoran Liang, Haiwei Dong, and Nikolaos Mavridis. Development of a novel conversational calculator based on remote online computation. Lecture Notes in Computer Science, 7663:142–151, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34475-6_18/.

Hassan:2012:LFA

[144] Saima Hassan, Abbas Khosravi, Jafreezal Jaafar, and Samir B. Belhaouari. Load forecasting accuracy through combination of trimmed forecasts. Lecture Notes in Computer Science, 7663:152–159, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34475-6_19/.

Szymanski:2012:SOM

[145] Julian Szymański and Włodzisław Duch. Self organizing maps for visualization of categories. Lecture Notes in Computer Science, 7663:

160–167, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34475-6_20/.

Anonymous:2012:FMg

[146] Anonymous. Front matter. Lecture Notes in Computer Science, 7663: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-34475-6/1.

Mezzoudj:2012:OMS

[147] Freha Mezzoudj and Assia Benyettou. On the optimization of multiclass support vector machines dedicated to speech recognition. Lecture Notes in Computer Science, 7664:1–8, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34481-7_1/.

Guo:2012:TSP

[148] Yangming Guo, Xiaolei Li, Guanghan Bai, and Jiezhong Ma. Time series prediction method based on LS-SVR with modified Gaussian RBF. Lecture Notes in Computer Science, 7664:9–17, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-34481-7_2/.

Wang:2012:IMI

[149] Ran Wang, Guoyou Wang, Zhong Chen, Jianguo Liu, and Yu Shi. An improved method of identification based on thermal palm vein image. Lecture Notes in Computer Science, 7664:18–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34481-7_3/.

Safa:2012:MEU

[150] Majeed Safa and Sandhya Samarasinghe. Modelling energy use and fuel consumption in wheat production using indirect factors and artificial neural networks. Lecture Notes in Computer Science, 7664:25– 32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34481-7_4/.

Tariq:2012:GMM

[151] Anam Tariq, Arslan Shaukat, and Shoab A. Khan. A Gaussian mixture model based system for detection of macula in fundus images. Lecture Notes in Computer Science, 7664:33–40, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34481-7_5/.

Brosch:2012:BSP

[152] Tobias Brosch and Heiko Neumann. The Brain’s sequential parallelism: Perceptual decision-making and early sensory responses. Lecture Notes in Computer Science, 7664:41–50, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34481-7_6/.

Koyama:2012:ENF

[153] Shinsuke Koyama. Estimating neural firing rates: An empirical Bayes approach. Lecture Notes in Computer Science, 7664:51–59, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34481-7_7/.

Sanchez:2012:AFR

[154] Jose Salvador Sánchez, Vicente García, and Ana Isabel Marqués. Assessment of financial risk prediction models with multi-criteria decision making methods. Lecture Notes in Computer Science, 7664:60– 67, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34481-7_8/.

Garcia:2012:IRP

[155] Vicente García, Ana Isabel Marqués, and Jose Salvador Sánchez. Improving risk predictions by preprocessing imbalanced credit data. Lecture Notes in Computer Science, 7664:68–75, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34481-7_9/.

Slim:2012:HVH

[156] Mohamed Aymen Slim, Maroua El Kastouri, Afef Abdelkrim, and Mohamed Benrejeb. Hybrid validation of handwriting process modelling. Lecture Notes in Computer Science, 7664:76–84, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-34481-7_10/.

Alalshekmubarak:2012:LHA

[157] Abdulrahman Alalshekmubarak, Amir Hussain, and Qiu-Feng Wang. Off-line handwritten Arabic word recognition using SVMs with normalized poly kernel. Lecture Notes in Computer Science, 7664:85–91, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34481-7_11/.

Xu:2012:BMV

[158] Jinhua Xu. Bayesian modeling of visual attention. Lecture Notes in Computer Science, 7664:92–99, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34481-7_12/.

Yin:2012:CEU

[159] Xu-Cheng Yin, Kaizhu Huang, Hong-Wei Hao, Khalid Iqbal, and Zhi-Bin Wang. Classifier ensemble using a heuristic learning with sparsity and diversity. Lecture Notes in Computer Science, 7664:100–107, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34481-7_13/.

Mitsuishi:2012:OSF

[160] Takashi Mitsuishi, Takanori Terashima, and Yasunari Shidama. Optimization of SIRMs fuzzy model using Łukasiewicz logic. Lecture Notes in Computer Science, 7664:108–116, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34481-7_14/.

Mahdavi:2012:SHL

[161] Nariman Mahdavi and Jürgen Kurths. Synchronization of Hopfield like chaotic neural networks with structure based learning. Lecture Notes in Computer Science, 7664:117–124, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34481-7_15/.

Paplinski:2012:ISO

[162] Andrew P. Papliński. Incremental self-organizing map (iSOM) in categorization of visual objects. Lecture Notes in Computer Science, 7664: 125–132, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34481-7_16/.

Cohen:2012:ENV

[163] Dror Cohen and Andrew P. Papliński. The elastic net as visual category representation: Visualisation and classification. Lecture Notes in Computer Science, 7664:133–140, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34481-7_17/.

Naze:2012:CMD

[164] Sebastien Naze and Jan Treur. A computational model for development of post-traumatic stress disorders by Hebbian learning. Lecture Notes in Computer Science, 7664:141–151, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34481-7_18/.

Samura:2012:PLS

[165] Toshikazu Samura, Yasuomi D. Sato, Yuji Ikegaya, Hatsuo Hayashi, and Takeshi Aihara. Power-law scaling of synchronization robustly reproduced in the hippocampal CA3 Slice culture model with small-world topology. Lecture Notes in Computer Science, 7664:152–159, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34481-7_19/.

Jamalabadi:2012:DBI

[166] Hamidreza Jamalabadi, Hossein Nasrollahi, Majid Nili Ahmadabadi, and Babak Nadjar Araabi. A dynamic bio-inspired model of categorization. Lecture Notes in Computer Science, 7664:160–167, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34481-7_20/.

Anonymous:2012:FMh

[167] Anonymous. Front matter. Lecture Notes in Computer Science, 7664: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-34481-7/1.

Sang:2012:CNN

[168] Do-Thanh Sang, Dong-Min Woo, and Dong-Chul Park. Centroid neural network with simulated annealing and its application to color image segmentation. Lecture Notes in Computer Science, 7665:1–8, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34487-9_1/.

Awad:2012:ENL

[169] Ali Awad. Efficient non-linear filter for impulse noise removal in document images. Lecture Notes in Computer Science, 7665:9–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34487-9_2/.

Bouaziz:2012:EFB

[170] Souhir Bouaziz, Habib Dhahri, and Adel M. Alimi. Evolving flexible Beta operator neural trees (FBONT) for time series forecasting. Lecture Notes in Computer Science, 7665:17–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34487-9_3/.

Fan:2012:NLF

[171] Wentao Fan and Nizar Bouguila. Nonparametric localized feature selection via a Dirichlet process mixture of generalized Dirichlet distributions. Lecture Notes in Computer Science, 7665:25–33, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34487-9_4/.

Treerattanapitak:2012:GAF

[172] Kiatichai Treerattanapitak and Chuleerat Jaruskulchai. Generalized agglomerative fuzzy clustering. Lecture Notes in Computer Science, 7665: 34–41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34487-9_5/.

Liu:2012:NSA

[173] Ming Liu, Lei Lin, Lili Shan, and Chengjie Sun. A novel self-adaptive clustering algorithm for dynamic data. Lecture Notes in Computer Science, 7665:42–49, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34487-9_6/.

Li:2012:ISS

[174] Chaojie Li, David Yang Gao, and Chao Liu. Impulsive synchronization of state delayed discrete complex networks with switching topology. Lecture Notes in Computer Science, 7665:50–57, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34487-9_7/.

Yuan:2012:SIS

[175] Liming Yuan, Songbo Liu, Qingcheng Huang, Jiafeng Liu, and Xianglong Tang. Salient instance selection for multiple-instance learning. Lecture Notes in Computer Science, 7665:58–67, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34487-9_8/.

Qingfeng:2012:MRM

[176] Meng Qingfeng, Du Jianguo, and Li Zhen. Motivating retail marketing efforts under fairness concerns in small-world networks: a multi-agent simulation. Lecture Notes in Computer Science, 7665:68–75, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34487-9_9/.

Gong:2012:AVG

[177] Dunwei Gong, Jian Chen, Xiaoyan Sun, and Yong Zhang. Application of variational granularity language sets in interactive genetic algorithms. Lecture Notes in Computer Science, 7665:76–83, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34487-9_10/.

Yokota:2012:LPC

[178] Tatsuya Yokota, Andrzej Cichocki, and Yukihiko Yamashita. Linked PARAFAC/CP tensor decomposition and its fast implementation for multi-block tensor analysis. Lecture Notes in Computer Science, 7665: 84–91, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34487-9_11/.

Xie:2012:HSI

[179] Gang Xie, Hong-Bo Guo, Yu-Chu Tian, and Maolin Tang. A humansimulated immune evolutionary computation approach. Lecture Notes in Computer Science, 7665:92–99, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34487-9_12/.

Zhenwei:2012:SBM

[180] Lu Zhenwei, Zhao Lingling, Su Xiaohong, and Ma Peijun. A STPHDbased multi-sensor fusion method. Lecture Notes in Computer Science, 7665:100–107, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-34487-9_13/.

Hara:2012:GSI

[181] Satoshi Hara and Takashi Washio. Group sparse inverse covariance selection with a dual augmented Lagrangian method. Lecture Notes in Computer Science, 7665:108–115, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34487-9_14/.

Liu:2012:MOL

[182] Yinglu Liu, Xu-Yao Zhang, Kaizhu Huang, Xinwen Hou, and Cheng-Lin Liu. Multiple outlooks learning with support vector machines. Lecture Notes in Computer Science, 7665:116–124, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34487-9_15/.

Srijith:2012:MTL

[183] P. K. Srijith and Shirish Shevade. Multi-task learning using shared and task specific information. Lecture Notes in Computer Science, 7665: 125–132, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34487-9_16/.

Eghbali:2012:LAF

[184] Sepehr Eghbali, Majid Nili Ahmadabadi, Babak Nadjar Araabi, and Maryam Mirian. Learning attentive fusion of multiple Bayesian network classifiers. Lecture Notes in Computer Science, 7665:133–140, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34487-9_17/.

Talaat:2012:MPL

[185] Amira Samy Talaat, Amir F. Atiya, Sahar A. Mokhtar, Ahmed Al-Ani, and Magda Fayek. Multiclass penalized likelihood pattern classification algorithm. Lecture Notes in Computer Science, 7665:141–148, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34487-9_18/.

Plonski:2012:SOM

[186] Piotr Płoński and Krzysztof Zaremba. Self-organising maps for classification with Metropolis–Hastings algorithm for supervision. Lecture Notes in Computer Science, 7665:149–156, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34487-9_19/.

Han:2012:DDD

[187] Yinghua Han, Jinkuan Wang, Qiang Zhao, and Peng Han. Decoupled 2-D DOA estimation algorithm based on cross-correlation matrix for coherently distributed source. Lecture Notes in Computer Science, 7665: 157–165, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34487-9_20/.

Anonymous:2012:FMi

[188] Anonymous. Front matter. Lecture Notes in Computer Science, 7665: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-34487-9/1.

Popescu:2012:NNL

[189] Theodor Dan Popescu. Neural network learning for blind source separation with application in dam safety monitoring. Lecture Notes in Computer Science, 7666:1–8, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34478-7_1/.

Mathews:2012:IBA

[190] Jayamol Mathews, Madhu S. Nair, and Liza Jo. Improved BTC algorithm for gray scale images using K-means quad clustering. Lecture Notes in Computer Science, 7666:9–17, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34478-7_2/.

Sulzbachner:2012:ONN

[191] Christoph Sulzbachner, Martin Humenberger, Ágoston Srp, and Ferenc Vajda. Optimization of a neural network for computer vision based fall detection with fixed-point arithmetic. Lecture Notes in Computer Science, 7666:18–26, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34478-7_3/.

Tu:2012:CRM

[192] Yan Tu and Zijiang Yang. Customer relationship management using partial focus feature reduction. Lecture Notes in Computer Science, 7666: 27–35, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34478-7_4/.

Akram:2012:ASG

[193] Muhammad Usman Akram, Mahmood Akhtar, and M. Younus Javed. An automated system for the grading of diabetic maculopathy in fundus images. Lecture Notes in Computer Science, 7666:36–43, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34478-7_5/.

Zou:2012:IMC

[194] Yun zhi Zou, Xin kun Wu, Wen bin Zhang, and Chang yin Sun. An iterative method for a class of generalized global dynamical system involving fuzzy mappings in Hilbert spaces. Lecture Notes in Computer Science, 7666:44–51, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34478-7_6/.

Kajornrit:2012:EMP

[195] Jesada Kajornrit, Kok Wai Wong, and Chun Che Fung. Estimation of missing precipitation records using modular artificial neural networks. Lecture Notes in Computer Science, 7666:52–59, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-34478-7_7/.

Morad:2012:OEA

[196] Ameer H. Morad. Office employees authentication based on E-exam techniques. Lecture Notes in Computer Science, 7666:60–65, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34478-7_8/.

Nagesh:2012:OSD

[197] Karthik Nagesh and M. Narasimha Murty. Obtaining single document summaries using latent Dirichlet allocation. Lecture Notes in Computer Science, 7666:66–74, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34478-7_9/.

Loo:2012:ORU

[198] Chu-Kiong Loo and Ali Memariani. Object recognition using sparse representation of overcomplete dictionary. Lecture Notes in Computer Science, 7666:75–82, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34478-7_10/.

Zheng:2012:PCA

[199] Jia Zheng, Mengyao Zhu, Junwei He, and Xiaoqing Yu. PEAQ compatible audio quality estimation using computational auditory model. Lecture Notes in Computer Science, 7666:83–90, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34478-7_11/.

Dewan:2012:SOC

[200] Sagar Dewan and Srinivasa Chakravarthy. A system for offline character recognition using auto-encoder networks. Lecture Notes in Computer Science, 7666:91–99, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34478-7_12/.

Yan:2012:TBB

[201] Xiaoyan Yan, Bo Liang, Tao Ban, Shanqing Guo, and Liming Wang. TrafficS: a behavior-based network traffic classification benchmark system with traffic sampling functionality. Lecture Notes in Computer Science, 7666:100–107, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-34478-7_13/.

Doush:2012:HSM

[202] Iyad Abu Doush. Harmony search with multi-parent crossover for solving IEEE-CEC2011 competition problems. Lecture Notes in Computer Science, 7666:108–114, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34478-7_14/.

Kacalak:2012:NII

[203] Wojciech Kacalak and Maciej Majewski. New intelligent interactive automated systems for design of machine elements and assemblies. Lecture Notes in Computer Science, 7666:115–122, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34478-7_15/.

Fu:2012:RSN

[204] Xiao Fu, Jin Liu, Haopeng Wang, Bin Zhang, and Rui Gao. Rough sets and neural networks based aerial images segmentation method. Lecture Notes in Computer Science, 7666:123–131, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34478-7_16/.

Ren:2012:FCN

[205] Shen Ren, Michael Barlow, and Hussein A. Abbass. Frontal cortex neural activities shift cognitive resources away from facial activities in real-time problem solving. Lecture Notes in Computer Science, 7666:

132–139, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34478-7_17/.

Kim:2012:IRT

[206] Geon min Kim, Chang hyun Kim, and Soo young Lee. Implement realtime polyphonic pitch detection and feedback system for the melodic instrument player. Lecture Notes in Computer Science, 7666:140–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34478-7_18/.

Kamiura:2012:CIS

[207] Naotake Kamiura, Ayumu Saitoh, Teijiro Isokawa, Nobuyuki Matsui, and Hitoshi Tabuchi. Classification of interview sheets using self-organizing maps for determination of ophthalmic examinations. Lecture Notes in Computer Science, 7666:148–155, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34478-7_19/.

Widodo:2012:SBR

[208] Slamet Widodo, Tomoo Shiigi, Naing Min Than, Yuichi Ogawa, and Naoshi Kondo. Sound-based ranging system in greenhouse environment with multipath effect compensation using artificial neural network. Lecture Notes in Computer Science, 7666:156–163, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34478-7_20/.

Anonymous:2012:FMj

[209] Anonymous. Front matter. Lecture Notes in Computer Science, 7666: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-34478-7/1.

Xiao:2012:DSNb

[210] Lu Xiao and Ming Fan. Does social network always promote entrepreneurial intentions? Part II: Empirical analysis. Lecture Notes in Computer Science, 7667:1–8, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34500-5_1/.

Wang:2012:RSM

[211] Xuzhi Wang, Yangyang Jia, Xiang Feng, Shuai Yu, and Hengyong Jiang. Rasterization system for mobile device. Lecture Notes in Computer Science, 7667:9–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-34500-5_2/.

Wang:2012:SRA

[212] Xuzhi Wang, Wei Xiong, Xiang Feng, Shuai Yu, and Hengyong Jiang. Study on rasterization algorithm for graphics acceleration system. Lecture Notes in Computer Science, 7667:17–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34500-5_3/.

Ito:2012:SLS

[213] Yoshifusa Ito, Hiroyuki Izumi, and Cidambi Srinivasan. Simultaneous learning of several Bayesian and Mahalanobis discriminant functions by a neural network with memory nodes. Lecture Notes in Computer Science, 7667:25–33, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34500-5_4/.

Sluzek:2012:RKM

[214] Andrzej Śluzek, Mariusz Paradowski, and Duanduan Yang. Reinforcement of keypoint matching by co-segmentation in object retrieval: Face recognition case study. Lecture Notes in Computer Science, 7667:34– 41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34500-5_5/.

Kudo:2012:ELG

[215] Hiroaki Kudo, Takuya Kume, and Noboru Ohnishi. Effect of luminance gradients in measurement of differential limen. Lecture Notes in Computer Science, 7667:42–49, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34500-5_6/.

Huang:2012:RCF

[216] Weicheng Huang, Shuichi Kurogi, and Takeshi Nishida. Robust controller for flexible specifications using difference signals and competitive associative nets. Lecture Notes in Computer Science, 7667:50–58, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34500-5_7/.

Ono:2012:MPD

[217] Kohei Ono, Shuichi Kurogi, and Takeshi Nishida. Moments of predictive deviations for ensemble diversity measures to estimate the performance of time series prediction. Lecture Notes in Computer Science, 7667:59–66, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34500-5_8/.

Morita:2012:SOT

[218] Satoru Morita. Self-organized three dimensional feature extraction of MRI and CT. Lecture Notes in Computer Science, 7667:67–74, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34500-5_9/.

Yao:2012:OBC

[219] Yao Yao, Xin Xin, and Ping Guo. OMP or BP? A comparison study of image fusion based on joint sparse representation. Lecture Notes in Computer Science, 7667:75–82, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34500-5_10/.

Yan:2012:IAS

[220] P. M. Yan, Meng Yang, Hui Huang, and J. F. Li. An improved approach to super resolution based on PET imaging. Lecture Notes in Computer Science, 7667:83–90, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34500-5_11/.

Wang:2012:PAC

[221] Zhi-Bin Wang, Hong-Wei Hao, Yan Li, Xu-Cheng Yin, and Shu Tian. Pedestrian analysis and counting system with videos. Lecture Notes in Computer Science, 7667:91–99, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34500-5_12/.

Zhao:2012:VLP

[222] Ya Zhao and Xiaodong Gu. Vehicle license plate localization and license number recognition using unit-linking pulse coupled neural network. Lecture Notes in Computer Science, 7667:100–108, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34500-5_13/.

Zhou:2012:RHL

[223] Shenghui Zhou, Qing Liu, Jianming Guo, and Yuanyuan Jiang. ROI-HOG and LBP based human detection via shape part-templates matching. Lecture Notes in Computer Science, 7667:109–115, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34500-5_14/.

Cancelliere:2012:MPI

[224] Rossella Cancelliere, Mario Gai, Thierry Artières, and Patrick Gallinari. Matrix pseudoinversion for image neural processing. Lecture Notes in Computer Science, 7667:116–125, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34500-5_15/.

Mattner:2012:LSB

[225] Jan Mattner, Sascha Lange, and Martin Riedmiller. Learn to swing up and balance a real pole based on raw visual input data. Lecture Notes in Computer Science, 7667:126–133, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34500-5_16/.

Lo:2012:GBB

[226] Alan W. Y. Lo, Benben Liu, and Ray C. C. Cheung. GPU-based biclustering for neural information processing. Lecture Notes in Computer Science, 7667:134–141, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34500-5_17/.

Sima:2012:CIS

[227] Haifeng Sima, Lixiong Liu, and Ping Guo. Color image segmentation based on regional saliency. Lecture Notes in Computer Science, 7667: 142–150, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34500-5_18/.

Wong:2012:MCD

[228] Man To Wong, Xiangjian He, Hung Nguyen, and Wei-Chang Yeh. Mass classification in digitized mammograms using texture features and artificial neural network. Lecture Notes in Computer Science, 7667:151– 158, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34500-5_19/.

Hu:2012:IDA

[229] Zhongyi Hu, Qing Liu, Shenghui Zhou, Mingjing Huang, and Fei Teng. Image dehazing algorithm based on atmosphere scatters approximation model. Lecture Notes in Computer Science, 7667:159–168, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34500-5_20/.

Anonymous:2012:FMk

[230] Anonymous. Front matter. Lecture Notes in Computer Science, 7667: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-34500-5/1.

Vaikuntanathan:2012:HCE

[231] Vinod Vaikuntanathan. How to compute on encrypted data. Lecture Notes in Computer Science, 7668:1–15, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34931-7_1/.

Dunkelman:2012:MEK

[232] Orr Dunkelman. From multiple encryption to knapsacks — efficient dissection of composite problems. Lecture Notes in Computer Science, 7668:16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-34931-7_2.

Kleinjung:2012:UCD

[233] Thorsten Kleinjung, Arjen K. Lenstra, Dan Page, and Nigel P. Smart. Using the cloud to determine key strengths. Lecture Notes in Computer Science, 7668:17–39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34931-7_3/.

Maji:2012:UCC

[234] Hemanta K. Maji, Manoj Prabhakaran, and Mike Rosulek. A unified characterization of completeness and triviality for secure function evaluation. Lecture Notes in Computer Science, 7668:40–59, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34931-7_4/.

Faust:2012:NMF

[235] Sebastian Faust, Markulf Kohlweiss, Giorgia Azzurra Marson, and Daniele Venturi. On the non-malleability of the Fiat–Shamir transform. Lecture Notes in Computer Science, 7668:60–79, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34931-7_5/.

Maitra:2012:ALS

[236] Arpita Maitra and Goutam Paul. Another look at symmetric incoherent optimal eavesdropping against BB84. Lecture Notes in Computer Science, 7668:80–99, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34931-7_6/.

Ganesh:2012:LLL

[237] Chaya Ganesh, Vipul Goyal, and Satya Lokam. On-line/off-line leakage resilient secure computation protocols. Lecture Notes in Computer Science, 7668:100–119, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34931-7_7/.

Carlet:2012:LSO

[238] Claude Carlet, Jean-Luc Danger, Sylvain Guilley, and Houssem Maghrebi. Leakage squeezing of order two. Lecture Notes in Computer Science, 7668: 120–139, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34931-7_8/.

Clavier:2012:RST

[239] Christophe Clavier, Benoit Feix, Georges Gagnerot, Christophe Giraud, and Mylène Roussellet. ROSETTA for single trace analysis. Lecture Notes in Computer Science, 7668:140–155, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34931-7_9/.

Lamberger:2012:CAH

[240] Mario Lamberger, Florian Mendel, and Vincent Rijmen. Collision attack on the Hamsi-256 compression function. Lecture Notes in Computer Science, 7668:156–171, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34931-7_10/.

Kortelainen:2012:GIH

[241] Tuomas Kortelainen, Ari Vesanen, and Juha Kortelainen. Generalized iterated hash functions revisited: New complexity bounds for multicollision attacks. Lecture Notes in Computer Science, 7668:172–190, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34931-7_11/.

Banik:2012:DFA

[242] Subhadeep Banik, Subhamoy Maitra, and Santanu Sarkar. A differential fault attack on the grain family under reasonable assumptions. Lecture Notes in Computer Science, 7668:191–208, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34931-7_12/.

Berger:2012:CPR

[243] Thierry P. Berger and Marine Minier. Cryptanalysis of pseudo-random generators based on vectorial FCSRs. Lecture Notes in Computer Science, 7668:209–224, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-34931-7_13/.

Derbez:2012:FCK

[244] Patrick Derbez, Pierre-Alain Fouque, and Jérémy Jean. Faster chosenkey distinguishers on reduced-round AES. Lecture Notes in Computer Science, 7668:225–243, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34931-7_14/.

Lu:2012:HOM

[245] Jiqiang Lu, Yongzhuang Wei, Jongsung Kim, and Enes Pasalic. The higher-order meet-in-the-middle attack and its application to the Camellia block cipher. Lecture Notes in Computer Science, 7668:244–264, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34931-7_15/.

Sasaki:2012:DSW

[246] Yu Sasaki. Double-SP is weaker than single-SP: Rebound attacks on Feistel ciphers with several rounds. Lecture Notes in Computer Science, 7668: 265–282, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34931-7_16/.

Wu:2012:AST

[247] Shengbao Wu and Mingsheng Wang. Automatic search of truncated impossible differentials for word-oriented block ciphers. Lecture Notes in Computer Science, 7668:283–302, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34931-7_17/.

Kim:2012:HSP

[248] Jung Woo Kim, Jungjoo Seo, Jin Hong, Kunsoo Park, and Sung-Ryul Kim. High-speed parallel implementations of the rainbow method in a heterogeneous system. Lecture Notes in Computer Science, 7668: 303–316, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34931-7_18/.

Bernstein:2012:CSD

[249] Daniel J. Bernstein and Tanja Lange. Computing small discrete logarithms faster. Lecture Notes in Computer Science, 7668:317–338, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34931-7_19/.

vonMaurich:2012:ESB

[250] Ingo von Maurich and Tim Güneysu. Embedded syndrome-based hashing. Lecture Notes in Computer Science, 7668:339–357, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34931-7_20/.

Anonymous:2012:FMl

[251] Anonymous. Front matter. Lecture Notes in Computer Science, 7668: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-34931-7/1.

Li:2012:AIP

[252] Mi Li, Dongning Han, Shengfu Lu, Zheng Liu, and Ning Zhong. Associative information processing in parahippocampal place area (PPA): An fMRI study. Lecture Notes in Computer Science, 7670:1–9, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35139-6_1/.

Tian:2012:LAO

[253] Qingyuan Tian. Language acquisition is an optimized process balanced between holistic and compositive. Lecture Notes in Computer Science, 7670:10–17, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35139-6_2/.

Lou:2012:ORT

[254] Chin Ian Lou, Daria Migotina, Joao P. Rodrigues, Joao Semedo, Feng Wan, and Peng Un Mak. Object recognition test in peripheral vision: a study on the influence of object color, pattern and shape. Lecture Notes in Computer Science, 7670:18–26, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35139-6_3/.

Fan:2012:BLA

[255] Sanxia Fan, Xuyan Wang, Zhizhou Liao, Zhoujun Long, Haiyan Zhou, and Yulin Qin. Basic level advantage during information retrieval: An ERP study. Lecture Notes in Computer Science, 7670:27–37, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35139-6_4/.

Long:2012:SLD

[256] Zhoujun Long, Xuyan Wang, Xiangsheng Shen, Sanxia Fan, Haiyan Zhou, and Yulin Qin. A skill learning during heuristic problem solving: An fMRI study. Lecture Notes in Computer Science, 7670:38–49, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35139-6_5/.

Wang:2012:RCL

[257] Xuyan Wang, Zhoujun Long, Sanxia Fan, Weiyan Yu, Haiyan Zhou, and Yulin Qin. The role of category label in adults’ inductive reasoning. Lecture Notes in Computer Science, 7670:50–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.com/chapter/10.1007/978-3-642-35139-6_6/.

Ueyama:2012:EVF

[258] Yuki Ueyama and Eizo Miyashita. Estimation of visual feedback contribution to limb stiffness in visuomotor control. Lecture Notes in Computer Science, 7670:61–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35139-6_7/.

Zhou:2012:RAP

[259] Haiyan Zhou, Yukun Xiong, Zhoujun Long, Sanxia Fan, Xuyan Wang, Yulin Qin, and Ning Zhong. Rule acquisition in the proceeding of heuristic Sudoku solving. Lecture Notes in Computer Science, 7670:

73–84, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35139-6_8/.

Mohammed:2012:BNT

[260] Redwan Abdo A. Mohammed, Samah Abdulfatah Mohammed, and Lars Schwabe. BatGaze: a new tool to measure depth features at the center of gaze during free viewing. Lecture Notes in Computer Science, 7670:85–96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35139-6_9/.

Mohammed:2012:BIA

[261] Redwan Abdo A. Mohammed and Lars Schwabe. A brain informatics approach to explain the oblique effect via depth statistics. Lecture Notes in Computer Science, 7670:97–106, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35139-6_10/.

Aziz:2012:AMT

[262] Azizi Ab Aziz, Faudziah Ahmad, and Houzifa M. Hintaya. An agent model for temporal dynamics analysis of a person with chronic fatigue syndrome. Lecture Notes in Computer Science, 7670:107–118, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35139-6_11/.

Wang:2012:NHM

[263] Zhijiang Wang, Jiming Liu, Ning Zhong, Yulin Qin, Haiyan Zhou, Jian Yang, and Kuncheng Li. A naïve hypergraph model of brain networks. Lecture Notes in Computer Science, 7670:119–129, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-35139-6_12/.

Choi:2012:EIP

[264] Do Young Choi, Kun Chang Lee, and Seong Wook Chae. The effect of individual psychological characteristics on creativity revelation: Emphasis with psychological empowerment and intrinsic motivation. Lecture Notes in Computer Science, 7670:130–139, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35139-6_13/.

Ryjov:2012:MOI

[265] Alexander Ryjov. Modeling and optimization of information retrieval for perception-based information. Lecture Notes in Computer Science, 7670: 140–149, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35139-6_14/.

Chen:2012:SPD

[266] Yu-Wen Chen and Hsing Mei. Sleep physiological dynamics simulation with fuzzy set. Lecture Notes in Computer Science, 7670:150–161, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-35139-6_15/.

DellArciprete:2012:PBM

[267] Lorenzo Dell’Arciprete, Brian Murphy, and Fabio Massimo Zanzotto. Parallels between machine and brain decoding. Lecture Notes in Computer Science, 7670:162–174, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35139-6_16/.

Ramirez:2012:DEE

[268] Rafael Ramirez and Zacharias Vamvakousis. Detecting emotion from EEG signals using the emotive epoc device. Lecture Notes in Computer Science, 7670:175–184, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-35139-6_17/.

Zhu:2012:EVG

[269] Guohun Zhu, Yan Li, and Peng Paul Wen. An efficient visibility graph similarity algorithm and its application on sleep stages classification. Lecture Notes in Computer Science, 7670:185–195, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-35139-6_18/.

Chou:2012:ITB

[270] Chun-An Chou, Kittipat “Bot” Kampa, Sonya H. Mehta, and Rosalia F. Tungaraza. Information-theoretic based feature selection for multi-voxel pattern analysis of fMRI data. Lecture Notes in Computer Science, 7670: 196–208, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35139-6_19/.

Atyabi:2012:MES

[271] Adham Atyabi, Sean P. Fitzgibbon, and David M. W. Powers. Multiplication of EEG samples through replicating, biasing, and overlapping. Lecture Notes in Computer Science, 7670:209–219, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-35139-6_20/.

Anonymous:2012:FMm

[272] Anonymous. Front matter. Lecture Notes in Computer Science, 7670: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-35139-6/1.

Kruegel:2012:FMS

[273] Christopher Kruegel. Fighting malicious software. Lecture Notes in Computer Science, 7671:1–15, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35130-3_1/.

Amrutkar:2012:VMS

[274] Chaitrali Amrutkar, Kapil Singh, Arunabh Verma, and Patrick Traynor. VulnerableMe: Measuring systemic weaknesses in mobile browser security. Lecture Notes in Computer Science, 7671:16–34, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35130-3_2/.

Ojamaa:2012:SAN

[275] Andres Ojamaa and Karl Düüna. Security assessment of node.js platform. Lecture Notes in Computer Science, 7671:35–43, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35130-3_3/.

Naldurg:2012:FDA

[276] Prasad Naldurg. Foundations of dynamic access control. Lecture Notes in Computer Science, 7671:44–58, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35130-3_4/.

Bijon:2012:RAR

[277] Khalid Zaman Bijon, Ram Krishnan, and Ravi Sandhu. Risk-aware RBAC sessions. Lecture Notes in Computer Science, 7671:59–74, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35130-3_5/.

Maity:2012:CTM

[278] Soumya Maity and Soumya K. Ghosh. A cognitive trust model for access control framework in MANET. Lecture Notes in Computer Science, 7671:75–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35130-3_6/.

Toumi:2012:TOT

[279] Khalifa Toumi, César Andrés, and Ana Cavalli. Trust-orBAC: a trust access control model in multi-organization environments. Lecture Notes in Computer Science, 7671:89–103, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35130-3_7/.

Bolton:2012:PI

[280] Matthew L. Bolton, Celeste M. Wallace, and Lenore D. Zuck. On policies and intents. Lecture Notes in Computer Science, 7671:104–118, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-35130-3_8/.

Swinnen:2012:PRP

[281] Arne Swinnen, Raoul Strackx, Pieter Philippaerts, and Frank Piessens. ProtoLeaks: a reliable and protocol-independent network covert channel. Lecture Notes in Computer Science, 7671:119–133, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35130-3_9/.

Paul:2012:KSS

[282] Goutam Paul, Ian Davidson, Imon Mukherjee, and S. S. Ravi. Keyless steganography in spatial domain using energetic pixels. Lecture Notes in Computer Science, 7671:134–148, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35130-3_10/.

Naskar:2012:FIR

[283] Ruchira Naskar and Rajat Subhra Chakraborty. Fuzzy inference rule based reversible watermarking for digital images. Lecture Notes in Computer Science, 7671:149–163, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35130-3_11/.

Natarajan:2012:STM

[284] Vivek Natarajan, Yi Yang, and Sencun Zhu. Secure trust metadata management for mobile ad-hoc networks. Lecture Notes in Computer Science, 7671:164–180, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-35130-3_12/.

Matam:2012:DAW

[285] Rakesh Matam and Somanath Tripathy. Defence against wormhole attacks in wireless mesh networks. Lecture Notes in Computer Science, 7671: 181–193, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35130-3_13/.

Agrawal:2012:NKU

[286] Sarita Agrawal, Rodrigo Roman, Manik Lal Das, Anish Mathuria, and Javier Lopez. A novel key update protocol in mobile sensor networks. Lecture Notes in Computer Science, 7671:194–207, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-35130-3_14/.

Bhattacharya:2012:ACL

[287] Anup Kumar Bhattacharya, Abhijit Das, Dipanwita Roychoudhury, and Aravind Iyer. Autonomous certification with list-based revocation for secure V2V communication. Lecture Notes in Computer Science, 7671: 208–222, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35130-3_15/.

Li:2012:MIT

[288] Weihan Li, Brajendra Panda, and Qussai Yaseen. Mitigating insider threat on database integrity. Lecture Notes in Computer Science, 7671: 223–237, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35130-3_16/.

Howlader:2012:CRM

[289] Jaydeep Howlader, Jayanta Kar, and Ashis Kumar Mal. Coercion resistant MIX for electronic auction. Lecture Notes in Computer Science, 7671: 238–248, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35130-3_17/.

Chandra:2012:VMA

[290] Girish Chandra and Divakar Yadav. Verification of money atomicity in digital cash based payment system. Lecture Notes in Computer Science, 7671:249–264, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-35130-3_18/.

Peinado:2012:CAT

[291] Alberto Peinado and Amparo Fúster-Sabater. Cryptographic analysis of a type of sequence generators. Lecture Notes in Computer Science, 7671: 265–276, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35130-3_19/.

Maity:2012:CLD

[292] Soumyadev Maity and R. C. Hansdah. Certificate-less on-demand public key management (CLPKM) for self-organized MANETs. Lecture Notes in Computer Science, 7671:277–293, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35130-3_20/.

Anonymous:2012:FMn

[293] Anonymous. Front matter. Lecture Notes in Computer Science, 7671: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-35130-3/1.

Han:2012:MIA

[294] Fengling Han and Ron van Schyndel. M-identity and its authentication protocol for secure mobile commerce applications. Lecture Notes in Computer Science, 7672:1–10, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35362-8_1/.

Srinivasan:2012:SSS

[295] Avinash Srinivasan and Jie Wu. SafeCode — safeguarding security and privacy of user data on stolen iOS devices. Lecture Notes in Computer Science, 7672:11–20, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35362-8_2/.

Braga:2012:PAI

[296] Alexandre Braga, Rafael Cividanes, Ismael Ávila, and Claudia Tambascia. Protection aspects of iconic passwords on mobile devices. Lecture Notes in Computer Science, 7672:21–32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35362-8_3/.

Graa:2012:DCF

[297] Mariem Graa, Nora Cuppens-Boulahia, Frédéric Cuppens, and Ana Cavalli. Detecting control flow in Smartphones: Combining static and dynamic analyses. Lecture Notes in Computer Science, 7672:33–47, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-35362-8_4/.

Abawajy:2012:MTE

[298] Jemal Abawajy and Andrei Kelarev. A multi-tier ensemble construction of classifiers for phishing email detection and filtering. Lecture Notes in Computer Science, 7672:48–56, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35362-8_5/.

Feng:2012:CFT

[299] Yong Feng, Bo Wang, Fengling Han, Xinghuo Yu, and Zahir Tari. Chattering-free terminal sliding-mode observer for anomaly detection. Lecture Notes in Computer Science, 7672:57–65, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-35362-8_6/.

Watters:2012:DID

[300] Paul A. Watters and Nigel Phair. Detecting illicit drugs on social media using automated social media intelligence analysis (ASMIA). Lecture Notes in Computer Science, 7672:66–76, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-35362-8_7/.

Timpanaro:2012:ICA

[301] Juan Pablo Timpanaro, Isabelle Chrisment, and Olivier Festor. Improving content availability in the I2P anonymous file-sharing environment. Lecture Notes in Computer Science, 7672:77–92, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35362-8_8/.

Stegelmann:2012:VVG

[302] Mark Stegelmann and Dogan Kesdogan. V2GPriv: Vehicle-to-grid privacy in the Smart Grid. Lecture Notes in Computer Science, 7672: 93–107, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35362-8_9/.

Angermeier:2012:SAS

[303] Daniel Angermeier, Konstantin Böttinger, Andreas Ibing, and Dieter Schuster. A secure architecture for smart meter systems. Lecture Notes in Computer Science, 7672:108–122, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35362-8_10/.

Yang:2012:NIB

[304] Geng Yang, Qiang Zhou, Xiaolong Xu, Jian Xu, and Chunming Rong. A novel identity-based key management and encryption scheme for distributed system. Lecture Notes in Computer Science, 7672:123–138, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-35362-8_11/.

Barbhuiya:2012:ABA

[305] Ferdous A. Barbhuiya, Tonmoy Saikia, and Sukumar Nandi. An anomaly based approach for HID attack detection using keystroke dynamics. Lecture Notes in Computer Science, 7672:139–152, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-35362-8_12/.

Srinivasan:2012:RAP

[306] Avinash Srinivasan and Lashidhar Chennupati. Robust authentication of public access points using digital certificates — a novel approach. Lecture Notes in Computer Science, 7672:153–164, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-35362-8_13/.

Fei:2012:GTK

[307] Han Fei, Qin Jing, Zhao Huawei, and Hu Jiankun. A general transformation from KP-ABE to searchable encryption. Lecture Notes in Computer Science, 7672:165–178, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35362-8_14/.

Ferretti:2012:SSC

[308] Luca Ferretti, Michele Colajanni, and Mirco Marchetti. Supporting security and consistency for cloud database. Lecture Notes in Computer Science, 7672:179–193, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35362-8_15/.

Petrlic:2012:PRE

[309] Ronald Petrlic. Proxy re-encryption in a privacy-preserving cloud computing DRM scheme. Lecture Notes in Computer Science, 7672:194– 211, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35362-8_16/.

Akimoto:2012:CBV

[310] Satoru Akimoto, Yoshiaki Hori, and Kouichi Sakurai. Collaborative behavior visualization and its detection by observing Darknet traffic. Lecture Notes in Computer Science, 7672:212–226, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-35362-8_17/.

Valli:2012:SSS

[311] Craig Valli. SSH — somewhat secure host. Lecture Notes in Computer Science, 7672:227–235, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35362-8_18/.

Skandhakumar:2012:PAC

[312] Nimalaprakasan Skandhakumar, Farzad Salim, Jason Reid, and Ed Dawson. Physical access control administration using building information models. Lecture Notes in Computer Science, 7672:236–250, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35362-8_19/.

Yang:2012:MFB

[313] Yongming Yang, Zulong Zhang, Fengling Han, and Kunming Lin. Multiple factors based evaluation of fingerprint images quality. Lecture Notes in Computer Science, 7672:251–264, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35362-8_20/.

Anonymous:2012:FMo

[314] Anonymous. Front matter. Lecture Notes in Computer Science, 7672: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-35362-8/1.

Elsayed:2012:INI

[315] Saber M. Elsayed, Ruhul A. Sarker, and Daryl L. Essam. The influence of the number of initial feasible solutions on the performance of an evolutionary optimization algorithm. Lecture Notes in Computer Science, 7673:1–11, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34859-4_1/.

Tagawa:2012:CDE

[316] Kiyoharu Tagawa. Concurrent differential evolution based on generational model for multi-core CPUs. Lecture Notes in Computer Science, 7673:12–21, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34859-4_2/.

Fu:2012:FMB

[317] Wenlong Fu, Mark Johnston, and Mengjie Zhang. Figure of merit based fitness functions in genetic programming for edge detection. Lecture Notes in Computer Science, 7673:22–31, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34859-4_3/.

Asco:2012:EAC

[318] Amadeo Ascó, Jason A. D. Atkin, and Edmund K. Burke. An evolutionary algorithm for the over-constrained airport baggage sorting station assignment problem. Lecture Notes in Computer Science, 7673: 32–41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34859-4_4/.

Phan:2012:NPS

[319] Dung H. Phan and Junichi Suzuki. A non-parametric statistical dominance operator for noisy multiobjective optimization. Lecture Notes in Computer Science, 7673:42–51, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34859-4_5/.

Khater:2012:ENG

[320] Marwa Khater, Elham Salehi, and Robin Gras. The emergence of new genes in EcoSim and its effect on fitness. Lecture Notes in Computer Science, 7673:52–61, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34859-4_6/.

Davarynejad:2012:MDG

[321] Mohsen Davarynejad, Zary Forghany, and Jan van den Berg. Massdispersed gravitational search algorithm for gene regulatory network model parameter identification. Lecture Notes in Computer Science, 7673: 62–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34859-4_7/.

Zhang:2012:DBA

[322] Bin Zhang, Kamran Shafi, and Hussein A. Abbass. A density based approach to the access point layout smart distribution grid design optimization problem. Lecture Notes in Computer Science, 7673:73–82, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34859-4_8/.

Ellabaan:2012:MMV

[323] Mostafa Ellabaan, Xianshun Chen, and Nguyen Quang Huy. Multi-modal valley-adaptive memetic algorithm for efficient discovery of first-order saddle points. Lecture Notes in Computer Science, 7673:83–92, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34859-4_9/.

Ishibuchi:2012:EFR

[324] Hisao Ishibuchi, Masakazu Yamane, and Yusuke Nojima. Ensemble fuzzy rule-based classifier design by parallel distributed fuzzy GBML algorithms. Lecture Notes in Computer Science, 7673:93–103, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34859-4_10/.

Kafafy:2012:HIH

[325] Ahmed Kafafy, Ahmed Bounekkar, and Stéphane Bonnevay. HEMH2: An improved hybrid evolutionary metaheuristics for 0/1 multiobjective knapsack problems. Lecture Notes in Computer Science, 7673:104–116, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34859-4_11/.

Rana:2012:GRD

[326] Prashant Singh Rana, Harish Sharma, Mahua Bhattacharya, and Anupam Shukla. Guided reproduction in differential evolution. Lecture Notes in Computer Science, 7673:117–127, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34859-4_12/.

Benlic:2012:SBL

[327] Una Benlic and Jin-Kao Hao. A study of breakout local search for the minimum sum coloring problem. Lecture Notes in Computer Science, 7673:128–137, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-34859-4_13/.

Nakata:2012:XAA

[328] Masaya Nakata, Pier Luca Lanzi, and Keiki Takadama. XCS with adaptive action mapping. Lecture Notes in Computer Science, 7673: 138–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34859-4_14/.

Vu:2012:DDG

[329] Cuong C. Vu, Lam Thu Bui, and Hussein A. Abbass. DEAL: a directionguided evolutionary algorithm. Lecture Notes in Computer Science, 7673: 148–157, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34859-4_15/.

Rosenthal:2012:IMS

[330] Susanne Rosenthal, Nail El-Sourani, and Markus Borschbach. Introduction of a mutation specific fast non-dominated sorting GA evolved for biochemical optimizations. Lecture Notes in Computer Science, 7673: 158–167, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34859-4_16/.

Sayed:2012:UHD

[331] Eman Sayed, Daryl Essam, and Ruhul Sarker. Using hybrid dependency identification with a memetic algorithm for large scale optimization problems. Lecture Notes in Computer Science, 7673:168–177, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34859-4_17/.

Chandra:2012:ACC

[332] Rohitash Chandra, Mengjie Zhang, and Lifeng Peng. Application of cooperative convolution optimization for ^{13}C metabolic flux analysis: Simulation of isotopic labeling patterns based on tandem mass spectrometry measurements. Lecture Notes in Computer Science, 7673:178–187, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34859-4_18/.

Huan:2012:ETP

[333] Hoang Xuan Huan, Dong Do Duc, and Nguyen Manh Ha. An efficient two-phase ant colony optimization algorithm for the closest string problem. Lecture Notes in Computer Science, 7673:188–197, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34859-4_19/.

Shafi:2012:EIM

[334] Kamran Shafi, Kathryn E. Merrick, and Essam Debie. Evolution of intrinsic motives in multi-agent simulations. Lecture Notes in Computer Science, 7673:198–207, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34859-4_20/.

Anonymous:2012:FMp

[335] Anonymous. Front matter. Lecture Notes in Computer Science, 7673: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-34859-4/1.

Wang:2012:ILP

[336] Chao Wang, Yunhong Wang, and Zhaoxiang Zhang. Incremental learning of patch-based bag of facial words representation for online face recognition in videos. Lecture Notes in Computer Science, 7674:1–9, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34778-8_1/.

Borowiak:2012:EAQ

[337] Adam Borowiak, Ulrich Reiter, and U. Peter Svensson. Evaluation of audio quality requirements over extended periods of time using Long duration audiovisual content. Lecture Notes in Computer Science, 7674:

10–20, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34778-8_2/.

Tao:2012:HCG

[338] Liang Tao and Horace H. S. Ip. Hashing with Cauchy graph. Lecture Notes in Computer Science, 7674:21–32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34778-8_3/.

Phan:2012:MED

[339] Sang Phan, Thanh Duc Ngo, Vu Lam, Son Tran, Duy-Dinh Le, Duc Anh Duong, and Shin’ichi Satoh. Multimedia event detection using segmentbased approach for motion feature. Lecture Notes in Computer Science, 7674:33–44, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34778-8_4/.

Romberg:2012:RFB

[340] Stefan Romberg, Moritz August, Christian X. Ries, and Rainer Lienhart. Robust feature bundling. Lecture Notes in Computer Science, 7674:45–56, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34778-8_5/.

Liang:2012:CGS

[341] Yang Liang, Mingli Song, Jiajun Bu, and Chun Chen. Colorization for gray scale facial image by locality-constrained linear coding. Lecture Notes in Computer Science, 7674:57–67, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34778-8_6/.

Zhou:2012:NEC

[342] Xiaozhou Zhou and Pierre Boulanger. New eye contact correction using radial basis function for wide baseline videoconference system. Lecture Notes in Computer Science, 7674:68–79, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34778-8_7/.

Yu:2012:IHF

[343] Wenxin Yu, Weichen Wang, Zhengyan Guo, and Satoshi Goto. An integrated hole-filling algorithm for view synthesis. Lecture Notes in Computer Science, 7674:80–92, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34778-8_8/.

Zhu:2012:RTC

[344] Shenggao Zhu, Hugh Anderson, and Ye Wang. A real-time on-chip algorithm for IMU-based gait measurement. Lecture Notes in Computer Science, 7674:93–104, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34778-8_9/.

Zhu:2012:RPC

[345] Shenggao Zhu, Hugh Anderson, and Ye Wang. Reducing the power consumption of an IMU-based gait measurement system. Lecture Notes in Computer Science, 7674:105–116, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34778-8_10/.

Wang:2012:SBE

[346] Yuxia Wang, Yuan Zhang, Rui Lu, and Pamela C. Cosman. SSIM-based end-to-end distortion modeling for H.264 video coding. Lecture Notes in Computer Science, 7674:117–128, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34778-8_11/.

Sun:2012:VOH

[347] Lei Sun, Zhenyu Liu, and Takeshi Ikenaga. A videoconferencing-oriented hybrid-domain H.264/SVC to H.264/AVC spatial transcoder. Lecture Notes in Computer Science, 7674:129–141, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34778-8_12/.

Tang:2012:RNE

[348] Chongwu Tang, Xiaokang Yang, and Guangtao Zhai. Robust noise estimation based on noise injection. Lecture Notes in Computer Science, 7674: 142–152, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34778-8_13/.

Wu:2012:IQA

[349] Jinjian Wu, Fei Qi, and Guangming Shi. Image quality assessment based on improved structural SIMilarity. Lecture Notes in Computer Science, 7674:153–163, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-34778-8_14/.

Chen:2012:MHT

[350] Chun-Chi Chen, Wen-Hsiao Peng, and Shih-Chun Chou. Multi-hypothesis temporal prediction using template matching prediction and block motion compensation for high efficiency video coding. Lecture Notes in Computer Science, 7674:164–175, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34778-8_15/.

Lu:2012:LES

[351] Hong Lu, Kai Chen, Guobao Jiang, Renzhong Wei, and Xiangyang Xue. Leveraging exemplar and saliency model for image search reranking. Lecture Notes in Computer Science, 7674:176–185, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34778-8_16/.

Qin:2012:RTM

[352] Shuxin Qin, Xiaoyang Zhu, Haitao Yu, Shuiying Ge, Yiping Yang, and Yongshi Jiang. Real-time markerless hand gesture recognition with depth camera. Lecture Notes in Computer Science, 7674:186–197, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34778-8_17/.

Xia:2012:STE

[353] Zhaoqiang Xia, Jinye Peng, Xiaoyi Feng, and Jianping Fan. Social tag enrichment via automatic abstract tag refinement. Lecture Notes in Computer Science, 7674:198–209, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34778-8_18/.

Quivy:2012:LOS

[354] Charles-Henri Quivy and Itsuo Kumazawa. “… it’s orange and small, and white stripes… ”. Lecture Notes in Computer Science, 7674:210– 221, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34778-8_19/.

Shukla:2012:ANR

[355] Arpit Shukla, Karan Nathwani, and Rajesh M. Hegde. An adaptive non reference anchor array framework for distant speech recognition. Lecture Notes in Computer Science, 7674:222–231, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34778-8_20/.

Anonymous:2012:FMq

[356] Anonymous. Front matter. Lecture Notes in Computer Science, 7674: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-34778-8/1.

Zhu:2012:EEP

[357] Yadong Zhu, Yuanhai Xue, Jiafeng Guo, Yanyan Lan, Xueqi Cheng, and Xiaoming Yu. Exploring and exploiting proximity statistic for information retrieval model. Lecture Notes in Computer Science, 7675: 1–13, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35341-3_1/.

Ji:2012:CIL

[358] Zongcheng Ji, Fei Xu, and Bin Wang. A category-integrated language model for question retrieval in community question answering. Lecture Notes in Computer Science, 7675:14–25, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-35341-3_2/.

Sakai:2012:RDS

[359] Tetsuya Sakai, Zhicheng Dou, Ruihua Song, and Noriko Kando. The reusability of a diversified search test collection. Lecture Notes in Computer Science, 7675:26–38, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35341-3_3/.

Sakai:2012:OCO

[360] Tetsuya Sakai and Makoto P. Kato. One click one revisited: Enhancing evaluation based on information units. Lecture Notes in Computer Science, 7675:39–51, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35341-3_4/.

Yue:2012:CAT

[361] Zhen Yue, Shuguang Han, and Daqing He. A comparison of action transitions in individual and collaborative exploratory Web search. Lecture Notes in Computer Science, 7675:52–63, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-35341-3_5/.

Qiu:2012:QOK

[362] Minghui Qiu, Yaliang Li, and Jing Jiang. Query-oriented keyphrase extraction. Lecture Notes in Computer Science, 7675:64–75, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35341-3_6/.

Wang:2012:ULT

[363] Jinpeng Wang, Wayne Xin Zhao, Rui Yan, Haitian Wei, Jian-Yun Nie, and Xiaoming Li. Using lexical and thematic knowledge for name disambiguation. Lecture Notes in Computer Science, 7675:76–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35341-3_7/.

Li:2012:SPS

[364] Xueni Li, Shaowu Zhang, Liang Yang, and Hongfei Lin. Sells out or piles up? A sentiment autoregressive model for predicting sales performance. Lecture Notes in Computer Science, 7675:89–102, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35341-3_8/.

Qi:2012:HIS

[365] Zhenyu Qi, Kang Liu, and Jun Zhao. Are human-input seeds good enough for entity set expansion? Seeds rewriting by leveraging Wikipedia semantic knowledge. Lecture Notes in Computer Science, 7675:103–113, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-35341-3_9/.

Puurula:2012:CMM

[366] Antti Puurula. Combining modifications to multinomial naive Bayes for text classification. Lecture Notes in Computer Science, 7675:114– 125, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35341-3_10/.

Mizuno:2012:OIW

[367] Junta Mizuno, Eric Nichols, Yotaro Watanabe, and Kentaro Inui. Organizing information on the Web through agreement-conflict relation classification. Lecture Notes in Computer Science, 7675:126–137, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35341-3_11/.

Yoshida:2012:ETS

[368] Mitsuo Yoshida and Yuki Arase. Exploiting Twitter for spiking query classification. Lecture Notes in Computer Science, 7675:138–149, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-35341-3_12/.

Ren:2012:WYS

[369] Kejiang Ren, Shaowu Zhang, and Hongfei Lin. Where are you settling down: Geo-locating Twitter users based on tweets and social networks. Lecture Notes in Computer Science, 7675:150–161, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35341-3_13/.

Chun:2012:DIM

[370] Chang-Woo Chun, Jung-Tae Lee, Seung-Wook Lee, and Hae-Chang Rim. Detecting informative messages based on user history in Twitter. Lecture Notes in Computer Science, 7675:162–173, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-35341-3_14/.

Jiang:2012:SPH

[371] Changhao Jiang, Min Zhang, Bin Gao, and Tie-Yan Liu. A study on potential head advertisers in sponsored search. Lecture Notes in Computer Science, 7675:174–186, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35341-3_15/.

Chang:2012:FTM

[372] Tianshu Chang, Hongfei Lin, and Yuan Lin. Feature transformation method enhanced vandalism detection in Wikipedia. Lecture Notes in Computer Science, 7675:187–198, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35341-3_16/.

Ye:2012:PQB

[373] Junting Ye, Yanan Qian, and Qinghua Zheng. PLIDMiner: a quality based approach for researcher’s homepage discovery. Lecture Notes in Computer Science, 7675:199–210, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35341-3_17/.

Lu:2012:EEH

[374] Kai Lu, Guanyuan Zhang, Rui Li, Shuai Zhang, and Bin Wang. Exploiting and exploring hierarchical structure in music recommendation. Lecture Notes in Computer Science, 7675:211–225, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-35341-3_18/.

Kaliciak:2012:EMI

[375] Leszek Kaliciak, Ben Horsburgh, Dawei Song, Nirmalie Wiratunga, and Jeff Pan. Enhancing music information retrieval by incorporating imagebased local features. Lecture Notes in Computer Science, 7675:226– 237, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35341-3_19/.

Shi:2012:YAS

[376] Liang Shi and Bin Wang. Yet another sorting-based solution to the reassignment of document identifiers. Lecture Notes in Computer Science, 7675:238–249, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-35341-3_20/.

Anonymous:2012:FMr

[377] Anonymous. Front matter. Lecture Notes in Computer Science, 7675: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-35341-3/1.

Deb:2012:IMS

[378] Debayan Deb and Kalyanmoy Deb. Investigation of mutation schemes in real-parameter genetic algorithms. Lecture Notes in Computer Science, 7677:1–8, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35380-2_1/.

Gao:2012:DHS

[379] Kaizhou Z. Gao. Discrete harmony search algorithm for dynamic FJSSP in remanufacturing engineering. Lecture Notes in Computer Science, 7677:9–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35380-2_2/.

Sarkar:2012:MIT

[380] Soham Sarkar and Swagatam Das. Multilevel image thresholding based on Tsallis entropy and differential evolution. Lecture Notes in Computer Science, 7677:17–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35380-2_3/.

Maity:2012:CBE

[381] Dipankar Maity and Udit Halder. Convergence and boundary estimation of the particle dynamics in generalized particle swarm optimization. Lecture Notes in Computer Science, 7677:25–32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35380-2_4/.

Das:2012:ADE

[382] Sudipta Das and Durbadal Mandal. Application of differential evolution with best of random mutation strategy on asymmetric location only synthesis of broadside circular antenna array. Lecture Notes in Computer Science, 7677:33–40, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35380-2_5/.

Varghese:2012:PEB

[383] Tinu Varghese and R. Sheela Kumari. Performance evaluation of bacterial foraging optimization algorithm for the early diagnosis and tracking of Alzheimer’s disease. Lecture Notes in Computer Science, 7677: 41–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35380-2_6/.

Senthilnath:2012:MSS

[384] J. Senthilnath and P. B. Shreyas. Multi-sensor satellite image analysis using niche genetic algorithm for flood assessment. Lecture Notes in Computer Science, 7677:49–56, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35380-2_7/.

Rathinam:2012:SEL

[385] Ananthanaryanan Rathinam and Ripunjoy Phukan. Solution to economic load dispatch problem based on FIREFLY algorithm and its comparison with BFO, CBFO-S and CBFO-hybrid. Lecture Notes in Computer Science, 7677:57–65, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35380-2_8/.

Ali:2012:MDM

[386] Layak Ali, Samrat L. Sabat, and Siba K. Udgata. MESFET DC model parameter extraction using adaptive accelerated exploration particle swarm optimizer. Lecture Notes in Computer Science, 7677:66–73, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35380-2_9/.

Reddy:2012:KGM

[387] Kalam Narendar Reddy and Vadlamani Ravi. Kernel group method of data handling: Application to regression problems. Lecture Notes in Computer Science, 7677:74–81, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35380-2_10/.

Agrawal:2012:EAG

[388] Sanjay Agrawal and Rutuparna Panda. An efficient algorithm for gray level image enhancement using cuckoo search. Lecture Notes in Computer Science, 7677:82–89, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35380-2_11/.

Panda:2012:MOW

[389] Sumanta Panda, Debadutta Mishra, and B. B. Biswal. A multi-objective workspace optimization of 3R manipulator using modified PSO. Lecture Notes in Computer Science, 7677:90–97, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-35380-2_12/.

Silva:2012:AGA

[390] Éderson R. Silva and Paulo R. Guardieiro. An analysis of genetic algorithm based anycast routing in delay and disruption tolerant networks. Lecture Notes in Computer Science, 7677:98–105, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35380-2_13/.

Bhattacharya:2012:RPO

[391] Bidishna Bhattacharya and Kamal Krishna Mandal. Reactive power optimization using hybrid cultural algorithm. Lecture Notes in Computer Science, 7677:106–115, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35380-2_14/.

Tudu:2012:TEF

[392] Bhimsen Tudu, Preetam Roy, and Sajjan Kumar. Techno-economic feasibility analysis of hybrid renewable energy system using improved version of particle swarm optimization. Lecture Notes in Computer Science, 7677: 116–123, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35380-2_15/.

Benala:2012:SEP

[393] Tirimula Rao Benala and Rajib Mall. Software effort prediction using fuzzy clustering and functional link artificial neural networks. Lecture Notes in Computer Science, 7677:124–132, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-35380-2_16/.

Nayak:2012:OPS

[394] Manas R. Nayak and Subrat K. Dash. Optimal placement and sizing of distributed generation in radial distribution system using differential evolution algorithm. Lecture Notes in Computer Science, 7677: 133–142, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35380-2_17/.

Ferrari:2012:CAR

[395] Daniel G. Ferrari and Leandro Nunes de Castro. Clustering algorithm recommendation: a meta-learning approach. Lecture Notes in Computer Science, 7677:143–150, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35380-2_18/.

Biswas:2012:CPB

[396] Subhodip Biswas, Digbalay Bose, and Souvik Kundu. A clustering particle based artificial bee colony algorithm for dynamic environment. Lecture Notes in Computer Science, 7677:151–159, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-35380-2_19/.

Kundu:2012:STL

[397] Souvik Kundu, Subhodip Biswas, and Swagatam Das. A selective teaching-learning based niching technique with local diversification strategy. Lecture Notes in Computer Science, 7677:160–168, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35380-2_20/.

Anonymous:2012:FMs

[398] Anonymous. Front matter. Lecture Notes in Computer Science, 7677: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-35380-2/1.

Sengamedu:2012:SAA

[399] Srinivasan H. Sengamedu. Scalable analytics — algorithms and systems. Lecture Notes in Computer Science, 7678:1–7, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.com/chapter/10.1007/978-3-642-35542-4_1/.

Agneeswaran:2012:BDT

[400] Vijay Srinivas Agneeswaran. Big-Data — theoretical, engineering and analytics perspective. Lecture Notes in Computer Science, 7678:8–15, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-35542-4_2/.

Basil:2012:CSM

[401] Tony Basil, Bollepalli S. Chandra, and Choudur Lakshminarayan. A comparison of statistical machine learning methods in heartbeat detection and classification. Lecture Notes in Computer Science, 7678: 16–25, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35542-4_3/.

Yasir:2012:EQO

[402] Ammar Yasir, Mittapally Kumara Swamy, Polepalli Krishna Reddy, and Subhash Bhalla. Enhanced query-by-object approach for information requirement elicitation in large databases. Lecture Notes in Computer Science, 7678:26–41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-35542-4_4/.

Gupta:2012:CCB

[403] Rajeev Gupta, Himanshu Gupta, and Mukesh Mohania. Cloud computing and big data analytics: What is new from databases perspective? Lecture Notes in Computer Science, 7678:42–61, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-35542-4_5/.

Reddy:2012:MVC

[404] Polepalli Krishna Reddy, Basi Bhaskar Reddy, and D. Rama Rao. A model of virtual crop labs as a cloud computing application for enhancing practical agricultural education. Lecture Notes in Computer Science, 7678:62–76, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35542-4_6/.

Yasir:2012:ESD

[405] Ammar Yasir, Mittapally Kumara Swamy, and Polepalli Krishna Reddy. Exploiting schema and documentation for summarizing relational databases. Lecture Notes in Computer Science, 7678:77–90, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35542-4_7/.

Nambiar:2012:FBS

[406] Ullas Nambiar, Tanveer Faruquie, Shamanth Kumar, Fred Morstatter, and Huan Liu. Faceted browsing over social media. Lecture Notes in Computer Science, 7678:91–100, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35542-4_8/.

Gupta:2012:ATE

[407] Anand Gupta, Manpreet Kathuria, Arjun Singh, Ashish Sachdeva, and Shruti Bhati. Analog textual entailment and spectral clustering (ATESC) based summarization. Lecture Notes in Computer Science, 7678:101– 110, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35542-4_9/.

Kumar:2012:EGP

[408] Jatin Kumar, Tushar Rao, and Saket Srivastava. Economics of gold price movement-forecasting analysis using macro-economic, investor fear and investor behavior features. Lecture Notes in Computer Science, 7678: 111–121, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35542-4_10/.

Ryu:2012:EMB

[409] Joung Woo Ryu, Mehmed M. Kantardzic, Myung-Won Kim, and A. Ra Khil. An efficient method of building an ensemble of classifiers in streaming data. Lecture Notes in Computer Science, 7678:122–133, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-35542-4_11/.

Mohanty:2012:EAB

[410] Sraban Kumar Mohanty and Gopalan Sajith. I/O efficient algorithms for block Hessenberg reduction using panel approach. Lecture Notes in Computer Science, 7678:134–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35542-4_12/.

Gupta:2012:LCT

[411] Anand Gupta, Prashant Khurana, and Raveena Mathur. Luring conditions and their proof of necessity through mathematical modelling. Lecture Notes in Computer Science, 7678:148–157, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-35542-4_13/.

Kim:2012:ERS

[412] Myung-Won Kim, Eun-Ju Kim, Won-Moon Song, Sung-Yeol Song, and A. Ra Khil. Efficient recommendation for smart TV contents. Lecture Notes in Computer Science, 7678:158–167, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-35542-4_14/.

Kumar:2012:MVS

[413] T. V. Vijay Kumar and Santosh Kumar. Materialized view selection using simulated annealing. Lecture Notes in Computer Science, 7678: 168–179, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35542-4_15/.

Anonymous:2012:BMb

[414] Anonymous. Back matter. Lecture Notes in Computer Science, 7678: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-35542-4/1.

Anonymous:2012:FMt

[415] Anonymous. Front matter. Lecture Notes in Computer Science, 7678: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-35542-4/1.

Morrisett:2012:SFM

[416] Greg Morrisett. Scalable formal machine models. Lecture Notes in Computer Science, 7679:1–3, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35308-6_1/.

Leroy:2012:MSC

[417] Xavier Leroy. Mechanized semantics for compiler verification. Lecture Notes in Computer Science, 7679:4–6, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-35308-6_2/.

Barthe:2012:ACA

[418] Gilles Barthe, Benjamin Grégoire, and César Kunz. Automation in computer-aided cryptography: Proofs, attacks and designs. Lecture Notes in Computer Science, 7679:7–8, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

accesspage/chapter/10.1007/978-3-642-35308-6_3.

Kobayashi:2012:PCH

[419] Naoki Kobayashi. Program certification by higher-order model checking. Lecture Notes in Computer Science, 7679:9–10, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/accesspage/chapter/10.1007/978-3-642-35308-6_4.

Robert:2012:FVA

[420] Valentin Robert and Xavier Leroy. A formally-verified alias analysis. Lecture Notes in Computer Science, 7679:11–26, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-35308-6_5/.

Zhao:2012:MVC

[421] Jianzhou Zhao and Steve Zdancewic. Mechanized verification of computing dominators for formalizing compilers. Lecture Notes in Computer Science, 7679:27–42, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35308-6_6/.

Mulligan:2012:COA

[422] Dominic P. Mulligan and Claudio Sacerdoti Coen. On the correctness of an optimising assembler for the Intel MCS-51 microprocessor. Lecture Notes in Computer Science, 7679:43–59, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-35308-6_7/.

Campbell:2012:ESC

[423] Brian Campbell. An executable semantics for CompCert C. Lecture Notes in Computer Science, 7679:60–75, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35308-6_8/.

Tollitte:2012:PCF

[424] Pierre-Nicolas Tollitte, David Delahaye, and Catherine Dubois. Producing certified functional code from inductive specifications. Lecture Notes in Computer Science, 7679:76–91, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35308-6_9/.

Bulwahn:2012:NQI

[425] Lukas Bulwahn. The new Quickcheck for Isabelle. Lecture Notes in Computer Science, 7679:92–108, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35308-6_10/.

Popescu:2012:PCN

[426] Andrei Popescu, Johannes Hölzl, and Tobias Nipkow. Proving concurrent noninterference. Lecture Notes in Computer Science, 7679:109– 125, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35308-6_11/.

Murray:2012:NOS

[427] Toby Murray, Daniel Matichuk, and Matthew Brassil. Noninterference for operating system kernels. Lecture Notes in Computer Science, 7679: 126–142, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35308-6_12/.

Vaynberg:2012:CVB

[428] Alexander Vaynberg and Zhong Shao. Compositional verification of a baby virtual memory manager. Lecture Notes in Computer Science, 7679: 143–159, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35308-6_13/.

Nakano:2012:SWJ

[429] Keisuke Nakano. Shall we juggle, coinductively? Lecture Notes in Computer Science, 7679:160–172, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35308-6_14/.

Accattoli:2012:PPA

[430] Beniamino Accattoli. Proof pearl: Abella formalization of λ-calculus cube property. Lecture Notes in Computer Science, 7679:173–187, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-35308-6_15/.

Chan:2012:SPP

[431] Hing-Lun Chan and Michael Norrish. A string of pearls: Proofs of Fermat’s Little Theorem. Lecture Notes in Computer Science, 7679: 188–207, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35308-6_16/.

Chaudhuri:2012:CPC

[432] Kaustuv Chaudhuri. Compact proof certificates for linear logic. Lecture Notes in Computer Science, 7679:208–223, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-35308-6_17/.

Doczkal:2012:CCM

[433] Christian Doczkal and Gert Smolka. Constructive completeness for modal logic with transitive closure. Lecture Notes in Computer Science, 7679: 224–239, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35308-6_18/.

Asperti:2012:RDE

[434] Andrea Asperti and Wilmer Ricciotti. Rating disambiguation errors. Lecture Notes in Computer Science, 7679:240–255, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-35308-6_19/.

Neron:2012:FPS

[435] Pierre Neron. A formal proof of square root and division elimination in embedded programs. Lecture Notes in Computer Science, 7679:

256–272, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35308-6_20/.

Anonymous:2012:FMu

[436] Anonymous. Front matter. Lecture Notes in Computer Science, 7679: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-35308-6/1.

Clarke:2012:MCS

[437] Edmund M. Clarke and William Klieber. Model checking and the state explosion problem. Lecture Notes in Computer Science, 7682: 1–30, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35746-6_1/.

Godefroid:2012:PLI

[438] Patrice Godefroid and Shuvendu K. Lahiri. From program to logic: An introduction. Lecture Notes in Computer Science, 7682:31–44, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-35746-6_2/.

Paulin-Mohring:2012:ICP

[439] Christine Paulin-Mohring. Introduction to the Coq proof-assistant for practical software verification. Lecture Notes in Computer Science, 7682: 45–95, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35746-6_3/.

Munoz:2012:ATP

[440] César A. Muñoz and Ramiro A. Demasi. Advanced theorem proving techniques in PVS and applications. Lecture Notes in Computer Science, 7682:96–132, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35746-6_4/.

Tschannen:2012:AVA

[441] Julian Tschannen and Carlo Alberto Furia. Automatic verification of advanced object-oriented features: The AutoProof approach. Lecture Notes in Computer Science, 7682:133–155, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35746-6_5/.

Herbert:2012:UDA

[442] Luke Herbert and K. Rustan M. Leino. Using Dafny, an automatic program verifier. Lecture Notes in Computer Science, 7682:156–181, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-35746-6_6/.

Anonymous:2012:BMc

[443] Anonymous. Back matter. Lecture Notes in Computer Science, 7682: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-35746-6/1.

Anonymous:2012:FMv

[444] Anonymous. Front matter. Lecture Notes in Computer Science, 7682: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-35746-6/1.

Gjoreski:2012:CBF

[445] Hristijan Gjoreski, Mitja Luštrek, and Matjaž Gams. Context-based fall detection using inertial and location sensors. Lecture Notes in Computer Science, 7683:1–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34898-3_1/.

Grosse-Puppendahl:2012:EAB

[446] Tobias Grosse-Puppendahl, Eugen Berlin, and Marko Borazio. Enhancing accelerometer-based activity recognition with capacitive proximity sensing. Lecture Notes in Computer Science, 7683:17–32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.com/chapter/10.1007/978-3-642-34898-3_2/.

Bongartz:2012:AUI

[447] Sara Bongartz, Yucheng Jin, Fabio Paternò, Joerg Rett, and Carmen Santoro. Adaptive user interfaces for smart environments with the support of model-based languages. Lecture Notes in Computer Science, 7683: 33–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34898-3_3/.

Osswald:2012:BSW

[448] Sebastian Osswald, Alexander Meschtscherjakov, Nicole Mirnig, and Karl-Armin Kraessig. Back of the steering wheel interaction: The car Braille keyer. Lecture Notes in Computer Science, 7683:49–64, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34898-3_4/.

Struse:2012:PCU

[449] Eric Struse, Julian Seifert, Sebastian Üllenbeck, Enrico Rukzio, and Christopher Wolf. PermissionWatcher: Creating user awareness of application permissions in mobile systems. Lecture Notes in Computer Science, 7683:65–80, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34898-3_5/.

Vaananen-Vainio-Mattila:2012:ENV

[450] Kaisa Väänänen-Vainio-Mattila, Tomi Haustola, Jonna Häkkilä, and Minna Karukka. Exploring non-verbal communication of presence between Young children and their parents through the embodied teddy bear. Lecture Notes in Computer Science, 7683:81–96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34898-3_6/.

Ijsselmuiden:2012:ABU

[451] Joris Ijsselmuiden, Ann-Kristin Grosselfinger, David Münch, and Michael Arens. Automatic behavior understanding in crisis response control rooms. Lecture Notes in Computer Science, 7683:97–112, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34898-3_7/.

Strasser:2012:CIE

[452] Ewald Strasser, Astrid Weiss, Thomas Grill, Sebastian Osswald, and Manfred Tscheligi. Combining implicit and explicit methods for the evaluation of an ambient persuasive factory display. Lecture Notes in Computer Science, 7683:113–128, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34898-3_8/.

Guivarch:2012:CAA

[453] Valérian Guivarch, Valérie Camps, and André Péninou. Context awareness in ambient systems by an adaptive multi-agent approach. Lecture Notes in Computer Science, 7683:129–144, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34898-3_9/.

Shell:2012:TFT

[454] Jethro Shell and Simon Coupland. Towards fuzzy transfer learning for intelligent environments. Lecture Notes in Computer Science, 7683: 145–160, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34898-3_10/.

Vatavu:2012:GPW

[455] Radu-Daniel Vatavu, Cătălin-Marian Chera, and Wei-Tek Tsai. Gesture profile for Web services: An event-driven architecture to support gestural interfaces for smart environments. Lecture Notes in Computer Science, 7683:161–176, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-34898-3_11/.

Chahuara:2012:UML

[456] Pedro Chahuara, Anthony Fleury, François Portet, and Michel Vacher. Using Markov logic network for on-line activity recognition from nonvisual home automation sensors. Lecture Notes in Computer Science, 7683:177–192, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-34898-3_12/.

Cvetkovic:2012:MCA

[457] Božidara Cvetković, Boštjan Kaluža, Mitja Luštrek, and Matjaž Gams. Multi-classifier adaptive training: Specialising an activity recognition classifier using semi-supervised learning. Lecture Notes in Computer Science, 7683:193–207, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-34898-3_13/.

Sehili:2012:SEA

[458] Mohamed A. Sehili, Benjamin Lecouteux, Michel Vacher, François Portet, and Dan Istrate. Sound environment analysis in smart home. Lecture Notes in Computer Science, 7683:208–223, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34898-3_14/.

Zachhuber:2012:CWO

[459] Doris Zachhuber, Thomas Grill, Ondrej Polacek, and Manfred Tscheligi. Contextual Wizard of Oz. Lecture Notes in Computer Science, 7683: 224–239, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34898-3_15/.

DeCarolis:2012:RUS

[460] Berardina De Carolis, Stefano Ferilli, and Nicole Novielli. Recognizing the user social attitude in multimodal interaction in smart environments. Lecture Notes in Computer Science, 7683:240–255, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34898-3_16/.

Shafti:2012:EFE

[461] Leila S. Shafti, Pablo A. Haya, Manuel García-Herranz, and Eduardo Pérez. Evolutionary feature extraction to infer behavioral patterns in ambient intelligence. Lecture Notes in Computer Science, 7683:256– 271, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34898-3_17/.

Kurdyukova:2012:PCP

[462] Ekaterina Kurdyukova, Stephan Hammer, and Elisabeth André. Personalization of content on public displays driven by the recognition of group context. Lecture Notes in Computer Science, 7683:272–287, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34898-3_18/.

Zaki:2012:TGA

[463] Michael Zaki and Peter Forbrig. Towards the generation of assistive user interfaces for smart meeting rooms based on activity patterns. Lecture Notes in Computer Science, 7683:288–295, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34898-3_19/.

Robben:2012:RDR

[464] Saskia Robben, Kyra Bergman, Sven Haitjema, Yannick de Lange, and Ben Kröse. Reducing dementia related wandering behaviour with an interactive wall. Lecture Notes in Computer Science, 7683:296–303, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34898-3_20/.

Anonymous:2012:FMw

[465] Anonymous. Front matter. Lecture Notes in Computer Science, 7683: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-34898-3/1.

Wang:2012:SLA

[466] Yanfeng Wang, Wenling Wu, Xiaoli Yu, and Lei Zhang. Security on LBlock against biclique cryptanalysis. Lecture Notes in Computer Science, 7690:1–14, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35416-8_1/.

Jia:2012:IID

[467] Keting Jia and Leibo Li. Improved impossible differential attacks on reduced-round MISTY1. Lecture Notes in Computer Science, 7690: 15–27, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35416-8_2/.

Tanaka:2012:EPE

[468] Satoshi Tanaka, Tung Chou, Bo-Yin Yang, Chen-Mou Cheng, and Kouichi Sakurai. Efficient parallel evaluation of multivariate quadratic polynomials on GPUs. Lecture Notes in Computer Science, 7690:28–42, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35416-8_3/.

Zhao:2012:EEV

[469] Wentao Zhao, Xin Hai, Shaojing Fu, Chao Li, and Yanfeng Wang. Enumeration of even-variable Boolean functions with maximum algebraic immunity. Lecture Notes in Computer Science, 7690:43–54, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35416-8_4/.

Seo:2012:MPM

[470] Hwajeong Seo and Howon Kim. Multi-precision multiplication for publickey cryptography on embedded microprocessors. Lecture Notes in Computer Science, 7690:55–67, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35416-8_5/.

Kim:2012:TPD

[471] Hyunmin Kim, Vladimir Rozic, and Ingrid Verbauwhede. Three phase dynamic current mode logic: a more secure DyCML to achieve a more balanced power consumption. Lecture Notes in Computer Science, 7690: 68–81, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35416-8_6/.

Park:2012:IDF

[472] JeaHoon Park and JaeCheol Ha. Improved differential fault analysis on block cipher ARIA. Lecture Notes in Computer Science, 7690: 82–95, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35416-8_7/.

Thorncharoensri:2012:MLC

[473] Pairat Thorncharoensri, Willy Susilo, and Yi Mu. Multi-level controlled signature. Lecture Notes in Computer Science, 7690:96–110, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-35416-8_8/.

Li:2012:TPC

[474] Liangze Li and Fan Zhang. Tate pairing computation on generalized Hessian curves. Lecture Notes in Computer Science, 7690:111– 123, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35416-8_9/.

Larangeira:2012:RCN

[475] Mario Larangeira and Keisuke Tanaka. Reduction-centric non-programmable security proof for the full domain hash in the random oracle model. Lecture Notes in Computer Science, 7690:124–143, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-35416-8_10/.

Kang:2012:AKM

[476] Hyun-Sun Kang and Chang-Seop Park. An authentication and key management scheme for the proxy mobile IPv6. Lecture Notes in Computer Science, 7690:144–160, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35416-8_11/.

Aszalos:2012:PAP

[477] László Aszalós and Andrea Huszti. Payment approval for PayWord. Lecture Notes in Computer Science, 7690:161–176, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-35416-8_12/.

Hwang:2012:ABA

[478] Jung Yeon Hwang, Sungwook Eom, Ku-Young Chang, Pil Joong Lee, and DaeHun Nyang. Anonymity-based authenticated key agreement with full binding property. Lecture Notes in Computer Science, 7690: 177–191, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35416-8_13/.

Oh:2012:SCW

[479] Junghoon Oh, Namheun Son, Sangjin Lee, and Kyungho Lee. A study for classification of Web browser log and timeline visualization. Lecture Notes in Computer Science, 7690:192–207, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-35416-8_14/.

Durvaux:2012:IPP

[480] François Durvaux, Benoît Gérard, Stéphanie Kerckhof, and François Koeune. Intellectual property protection for integrated systems using soft physical hash functions. Lecture Notes in Computer Science, 7690: 208–225, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35416-8_15/.

Liu:2012:VAD

[481] Shun-Te Liu, Yi-Ming Chen, and Hui-Ching Hung. N-victims: An approach to determine N-victims for APT investigations. Lecture Notes in Computer Science, 7690:226–240, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35416-8_16/.

Choi:2012:EFM

[482] Jaeun Choi, Gisung Kim, Tae Ghyoon Kim, and Sehun Kim. An efficient filtering method for detecting malicous Web pages. Lecture Notes in Computer Science, 7690:241–253, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35416-8_17/.

Lin:2012:LCS

[483] Eric Lin, John Aycock, and Mohammad Mannan. Lightweight client-side methods for detecting email forgery. Lecture Notes in Computer Science, 7690:254–269, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-35416-8_18/.

Choi:2012:ATB

[484] Yang seo Choi, Ik-Kyun Kim, Jin-Tae Oh, and Jong-Soo Jang. AIGG threshold based HTTP GET flooding attack detection. Lecture Notes in Computer Science, 7690:270–284, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35416-8_19/.

Kim:2012:IGW

[485] Hyunjoo Kim, Byoungkoo Kim, Daewon Kim, Ik-Kyun Kim, and Tai-Myoung Chung. Implementation of GESNIC for Web server protection against HTTP GET flooding attacks. Lecture Notes in Computer Science, 7690:285–295, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-35416-8_20/.

Anonymous:2012:FMx

[486] Anonymous. Front matter. Lecture Notes in Computer Science, 7690: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-35416-8/1.

Nguyen:2012:CCIa

[487] Ngoc-Thanh Nguyen, Kiem Hoang, and Piotr Jȩdrzejowicz, editors. Computational Collective Intelligence. Technologies and Applications: 4th International Conference, ICCCI 2012, Ho Chi Minh City, Vietnam, November 28–30, 2012, Proceedings, Part I, volume 7653 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34629-4 (print), 3-642-34630-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-34630-9.

Nguyen:2012:CCIb

[488] Ngoc-Thanh Nguyen, Kiem Hoang, and Piotr Jȩdrzejowicz, editors. Computational Collective Intelligence. Technologies and Applications: 4th International Conference, ICCCI 2012, Ho Chi Minh City, Vietnam, November 28–30, 2012, Proceedings, Part II, volume 7654 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34706-1 (print), 3-642-34707-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-34707-8.

Barneva:2012:CIA

[489] Reneta P. Barneva, Valentin E. Brimkov, and Jake K. Aggarwal, editors. Combinatorial Image Analaysis: 15th International Workshop, IW-CIA 2012, Austin, TX, USA, November 28–30, 2012. Proceedings, volume 7655 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34731-2 (print), 3-642-34732-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-642-
34732-0.

Wang:2012:ACA

[490] Xiaoyun Wang and Kazue Sako, editors. Advances in Cryptology — ASI-ACRYPT 2012: 18th International Conference on the Theory and Application of Cryptology and Information Security, Beijing, China, December 2–6, 2012. Proceedings, volume 7658 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34960-9 (print), 3-642-34961-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-34961-4.

Even:2012:DAA

[491] Guy Even and Dror Rawitz, editors. Design and Analysis of Algorithms: First Mediterranean Conference on Algorithms, MedAlg 2012, Kibbutz Ein Gedi, Israel, December 3–5, 2012. Proceedings, volume 7659 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34861-0 (print), 3-642-34862-9 (ebook). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL

http://www.springerlink.com/content/978-3-642-34862-4.

Chen:2012:FIS

[492] Li Chen, Alexander Felfernig, Jiming Liu, and Zbigniew W. Raś, editors. Foundations of Intelligent Systems: 20th International Symposium, ISMIS 2012, Macau, China, December 4–7, 2012. Proceedings, volume 7661 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34623-5 (print), 3-642-34624-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-34624-8.

Huang:2012:NIPa

[493] Tingwen Huang, Zhigang Zeng, Chuandong Li, and Chi Sing Leung, editors. Neural Information Processing: 19th International Conference, ICONIP 2012, Doha, Qatar, November 12–15, 2012, Proceedings, Part I, volume 7663 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34474-7 (print), 3-642-34475-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-34475-6.

Huang:2012:NIPb

[494] Tingwen Huang, Zhigang Zeng, Chuandong Li, and Chi Sing Leung, editors. Neural Information Processing: 19th International Conference, ICONIP 2012, Doha, Qatar, November 12–15, 2012, Proceedings, Part II, volume 7664 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34480-1 (print), 3-642-34481-X (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-34481-7.

Huang:2012:NIPc

[495] Tingwen Huang, Zhigang Zeng, Chuandong Li, and Chi Sing Leung, editors. Neural Information Processing: 19th International Conference, ICONIP 2012, Doha, Qatar, November 12–15, 2012, Proceedings, Part III, volume 7665 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34486-0 (print), 3-642-34487-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-34487-9.

Huang:2012:NIPd

[496] Tingwen Huang, Zhigang Zeng, Chuandong Li, and Chi Sing Leung, editors. Neural Information Processing: 19th International Conference, ICONIP 2012, Doha, Qatar, November 12–15, 2012, Proceedings, Part IV, volume 7666 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34477-1 (print), 3-642-34478-X (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-34478-7.

Huang:2012:NIPe

[497] Tingwen Huang, Zhigang Zeng, Chuandong Li, and Chi Sing Leung, editors. Neural Information Processing: 19th International Conference, ICONIP 2012, Doha, Qatar, November 12–15, 2012, Proceedings, Part V, volume 7667 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34499-2 (print), 3-642-34500-X (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-34500-5.

Galbraith:2012:PCI

[498] Steven Galbraith and Mridul Nandi, editors. Progress in Cryptology — INDOCRYPT 2012: 13th International Conference on Cryptology in India, Kolkata, India, December 9–12, 2012. Proceedings, volume 7668 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34930-7 (print), 3-642-34931-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-34931-7.

Zanzotto:2012:BII

[499] Fabio Massimo Zanzotto, Shusaku Tsumoto, Niels Taatgen, and Yiyu Yao, editors. Brain Informatics: International Conference, BI 2012, Macau, China, December 4–7, 2012. Proceedings, volume 7670 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-35138-7 (print), 3-642-35139-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-35139-6.

Venkatakrishnan:2012:ISS

[500] Venkat Venkatakrishnan and Diganta Goswami, editors. Information Systems Security: 8th International Conference, ICISS 2012, Guwahati, India, December 15–19, 2012. Proceedings, volume 7671 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-35129-8 (print), 3-642-35130-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-35130-3.

Xiang:2012:CSS

[501] Yang Xiang, Javier Lopez, C.-C. Jay Kuo, and Wanlei Zhou, editors. Cyberspace Safety and Security: 4th International Symposium, CSS 2012, Melbourne, Australia, December 12–13, 2012. Proceedings, volume 7672 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-35361-4 (print), 3-642-35362-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-35362-8.

Bui:2012:SEL

[502] Lam Thu Bui, Yew Soon Ong, Nguyen Xuan Hoai, Hisao Ishibuchi, and Ponnuthurai Nagaratnam Suganthan, editors. Simulated Evolution and Learning: 9th International Conference, SEAL 2012, Hanoi, Vietnam, December 16–19, 2012. Proceedings, volume 7673 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34858-0 (print), 3-642-34859-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-34859-4.

Lin:2012:AMI

[503] Weisi Lin, Dong Xu, Anthony Ho, Jianxin Wu, Ying He, Jianfei Cai, Mohan Kankanhalli, and Ming-Ting Sun, editors. Advances in Multimedia Information Processing — PCM 2012: 13th Pacific-Rim Conference on Multimedia, Singapore, December 4–6, 2012. Proceedings, volume 7674 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34777-0 (print), 3-642-34778-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-34778-8.

Hou:2012:IRT

[504] Yuexian Hou, Jian-Yun Nie, Le Sun, Bo Wang, and Peng Zhang, editors. Information Retrieval Technology: 8th Asia Information Retrieval Societies Conference, AIRS 2012, Tianjin, China, December 17–19, 2012. Proceedings, volume 7675 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-35340-1 (print), 3-642-35341-X (e-book). ISSN 0302-9743 (print), 1611-

3349 (electronic). LCCN ???? URL http://www.springerlink.com/
content/978-3-642-35341-3.

Panigrahi:2012:SEM

[505] Bijaya Ketan Panigrahi, Swagatam Das, Ponnuthurai Nagaratnam Suganthan, and Pradipta Kumar Nanda, editors. Swarm, Evolutionary, and Memetic Computing: Third International Conference, SEMCCO 2012, Bhubaneswar, India, December 20–22, 2012. Proceedings, volume 7677 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-35379-7 (print), 3-642-35380-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-35380-2.

Srinivasa:2012:BDA

[506] Srinath Srinivasa and Vasudha Bhatnagar, editors. Big Data Analytics: First International Conference, BDA 2012, New Delhi, India, December 24–26, 2012. Proceedings, volume 7678 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-35541-2 (print), 3-642-35542-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-35542-4.

Hawblitzel:2012:CPP

[507] Chris Hawblitzel and Dale Miller, editors. Certified Programs and Proofs: Second International Conference, CPP 2012, Kyoto, Japan, December 13– 15, 2012. Proceedings, volume 7679 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-35307-X (print), 3-642-35308-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-35308-6.

Meyer:2012:TPS

[508] Bertrand Meyer and Martin Nordio, editors. Tools for Practical Software Verification: LASER, International Summer School 2011, Elba Island, Italy, Revised Tutorial Lectures, volume 7682 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-35745-8 (print), 3-642-35746-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.

springerlink.com/content/978-3-642-35746-6.

Paterno:2012:AIT

[509] Fabio Paternò, Boris de Ruyter, Panos Markopoulos, Carmen Santoro, Evert van Loenen, and Kris Luyten, editors. Ambient Intelligence: Third International Joint Conference, AmI 2012, Pisa, Italy, November 13–15, 2012. Proceedings, volume 7683 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34897-1 (print), 3-642-34898-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-34898-3.

Lee:2012:ISA

[510] Dong Hoon Lee and Moti Yung, editors. Information Security Applications: 13th International Workshop, WISA 2012, Jeju Island, Korea, August 16–18, 2012, Revised Selected Papers, volume 7690 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-35415-7 (print), 3-642-35416-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-35416-8.